### v2

# Without Paired Labeled Data: An End-to-End Self-Supervised Paradigm for UAV-View Geo-Localization

Abstract
UAV-View Geo-Localization (UVGL) aims to achieve accurate localization of unmanned aerial vehicles (UAVs) by retrieving the most relevant GPS-tagged satellite images. However, existing methods heavily rely on pre-paired UAV-satellite images for supervised learning. Such dependency not only incurs high annotation costs but also severely limits scalability and practical deployment in open-world UVGL scenarios. To address these limitations, we propose an end-to-end self-supervised UVGL method. Our method leverages a shallow backbone network to extract initial features, employs clustering to generate pseudo labels, and adopts a dual-path contrastive learning architecture to learn discriminative intra-view representations. Furthermore, our method incorporates two core modules, the dynamic hierarchical memory learning module and the information consistency evolution learning module. The dynamic hierarchical memory learning module combines short-term and long-term memory to enhance intra-view feature consistency and discriminability. Meanwhile, the information consistency evolution learning module leverages a neighborhood-driven dynamic constraint mechanism to systematically capture implicit cross-view semantic correlations, thereby improving cross-view feature alignment. To further stabilize and strengthen the self-supervised training process, a pseudo-label enhancement strategy is introduced, which refines the quality of pseudo supervision. Our method ultimately constructs a unified cross-view feature representation space under self-supervised settings. Extensive experiments on three public benchmark datasets demonstrate that the proposed method consistently outperforms existing self-supervised methods and even surpasses several state-of-the-art supervised methods. These results validate the effectiveness and generalizability of our method, highlighting its potential for real-world deployment in UVGL. Our code is available at https://github.com/ISChenawei/DMNIL.
Index Terms:
UAV-View Geo-Localization, Self-Supervised Learning, Contrastive Learning.I Introduction
UAV-View Geo-Localization (UVGL) aims to achieve high-precision localization of UAV in GPS-denied scenarios by retrieving GPS-tagged satellite images that correspond to UAV-captured near-ground view images [1]. This capability is critical for applications such as autonomous navigation, disaster rescue, and emergency response. In recent years, existing UVGL methods have primarily relied on supervised learning using strictly paired UAV-satellite images [2, 3, 4]. Although these supervised methods have achieved encouraging success, they also reveal several inherent limitations: (1) acquiring instance-level paired UAV-satellite images is highly costly [5], leading to increased training overhead; (2) large amounts of unpaired cross-view image resources remain underutilized; and (3) when applied to new scenarios, these methods often require recollection and re-pairing UAV-satellite images for training to adapt to domain shifts. These limitations significantly hinder the scalability and practical deployment of UVGL in real-world scenarios. Thereby, it is necessary and of great significance to explore a self-supervised UVGL method that can effectively mine latent cross-view correlations from large-scale unpaired UAV-satellite images.
In recent years, self-supervised person re-identification (ReID) methods [6, 7] have achieved remarkable progress, offering crucial theoretical foundations for UVGL. Typically, These methods employ clustering algorithms [8, 9] on features extracted by a backbone network to generate pseudo-labels, which serve as the initial supervision signals for self-supervised learning. However, the quality of these pseudo-labels is heavily dependent on the discriminative capability of backbone network. As illustrated in Fig. 1, compared with conventional ReID, UVGL faces additional challenges. These factors make it more difficult for the backbone to extract discriminative features, as illustrated in Fig. 2. When pseudo-labels contain a high error rate and the model lacks the capacity to capture intra-view and cross-view consistency and discriminative features, the self-supervised learning process becomes prone to catastrophic negative optimization.
In this paper, we propose a novel end-to-end self-supervised UVGL method. A lightweight, shallow ConvNeXt-Tiny backbone network [10] is employed to extract features, which are subsequently clustered to generate pseudo-labels. Combined with a dual-path contrastive learning strategy [17, 18], these components form the baseline network of our self-supervised method for UVGL to learn intra-view feature representations. Building upon this baseline, we proposed the dynamic hierarchical memory learning (DHML) module and the information consistency evolution learning (ICEL) module as two key components of our method. The DHML module captures the relationship between short-term and long-term memory to learn the discriminability of intra-view feature representations. The ICEL module constructs a neighborhood-driven dynamic constraint mechanism based on feature distributions. By integrating neighborhood consistency constraints with mutual information optimization, it effectively aligns cross-view feature representations within a shared embedding space, promoting both consistency and discriminative learning. Additionally, to improve the reliability of pseudo-labels, we introduce a pseudo-label enhancement (PLE) strategy, which contributes to a more stable and effective self-supervised feature learning process. The main contributions can be summarized as follows:
We propose an end-to-end self-supervised method for UVGL. Built upon a shallow backbone and a dual-path contrastive learning strategy, our method integrates dynamic hierarchical memory learning and information consistency evolution learning to effectively capture and enhance feature consistency and discriminability for both intra-view and cross-view representation learning.
We propose a dynamic memory hierarchical learning module that captures the relationships among different levels of memory to adaptively model intra-class and inter-class feature dynamics within the intra-view, thereby enhancing the discriminability of feature representations.
we propose a novel information consistency evolution learning module which combines neighborhood consistency and mutual information optimization to learn consistent and discriminative cross-view representations.
Extensive experiments demonstrate that our method, utilizing only a lightweight backbone, achieves state-of-the-art performance on three UVGL benchmarks, outperforming even certain supervised methods.
The remainder of this paper is organized as follows. Section II systematically reviews prior research. In Section III, the proposed self-supervised UVGL method is presented in detail. The experimental results are reported and analyzed in Section IV. Finally, conclusions are outlined in Section V.
II RELATED WORKS
This section establishes a systematic and comprehensive theoretical framework to support the development of the proposed self-supervised UVGL method. Although this work focuses on UVGL as a sub-task of cross-view geo-localization (CVGL), we first review the overall progress in supervised CVGL. We then discuss recent advances in self-supervised CVGL methods. Finally, we summarize main methodes in self-supervised ReID.
II-A Supervised Cross-View Geo-localization
CVGL aims to achieve high-precision localization of query images by retrieving GPS-tagged satellite images. Early research in CVGL primarily focused on the ground-view CVGL [19, 20, 21], leading to the development of several benchmark datasets, such as CVUSA [22], CVACT [23], and VIGOR [24]. To address the challenge of ground-view CVGL, researchers have proposed various methods. CVM-Net [25] employed NetVLAD [26] to encode images as global descriptors, thereby reducing visual discrepancies caused by viewpoint variations. TransGeo [27] utilized attention-based zooming and sharpness-aware optimization to enhance detail learning and improve model generalization. GeoDTR [20] leveraged a Transformer-based [28] feature extractor to disentangle geometric features, effectively addressing perspective shifts. Sample4Geo [29] introduced a hard negative mining strategy, focusing on challenging negative samples to further improve discriminative capability.
With the continuous development and application of drone technology, UVGL has gradually become an important research direction and achieves significant progress [30, 31, 32]. Several benchmark datasets for UVGL have been proposed, such as University-1652[12], SUES-200, [33] and DenseUAV [34], which have driven further research in this field. In this scenario, a variety of methods have been proposed by researchers. LPN [35] mined fine-grained features through local pattern partitioning, improving the ability to capture local details. IFS [36] utilized a multi-branch strategy to effectively fuse global and local features, enhancing the diversity and robustness of feature representations. CAMP [30] improved cross-view matching accuracy by contrastive attribute mining and position-aware partitioning, while DAC [4] enhanced feature consistency and matching accuracy by introducing domain alignment and scene consistency constraints. MEAN [2] adopted progressive embedding diversification, global-to-local associations, and cross-domain enhanced alignment to further improve feature representation and alignment.
Nevertheless, the remarkable performances exhibited by these UVGL methods require extensive pre-processed UAV-satellite image pairs. In this work, we pivot our focus to the realm of self-supervised UVGL, where prior pairing relationships are absent, presenting important applications for real-world UVGL deployments.
II-B Self-Supervised Cross-View Geo-localization
Due to the dependence of CVGL on extensive pre-processed image pairs, which significantly increases costs and limits applications for real-world, researchers have begun to explore self-supervised CVGL methods. The work in [37] proposed a self-supervised method that leveraged correspondence-free projection to transform ground panorama images into bird’s-eye view images. It also applied CycleGAN [38] to generate fake satellite images, coupled with an additional SAFA module [39] to reduce discrepancies. However, this explicit alignment method has inherent limitations in capturing the deep semantic relationships of cross-view features and may introduce noise. Furthermore, the non-end-to-end training framework employed significantly increases training time and computational costs, reducing the efficiency and practical applicability. Additionally, existing work [5] utilized frozen foundation models (FMs) [40, 41] to extract features, and a self-supervised adapter was introduced to mitigate visual discrepancies between different point of view. However, this method heavily relied on the ability of FMs for feature extraction. When FMs fail to effectively handle viewpoint discrepancies, the performance of the adapter is also affected. To reduce reliance on the feature extractor, CDIKTNet [42] trained with a small amount of paired UAV-satellite images to optimize the initial feature distribution, and then performed clustering to generate pseudo-labels, thereby improving the quality of pseudo-labels generation. However, this method inherently depends on the feature extractor, as it still requires limited paired UAV-satellite images for effective learning.
The above methods have made meaningful exploratory attempts. However, these methods still face challenges, such as the additional computation time and overhead introduced by non-end-to-end frameworks, as well as dependence on the capability of feature extractor. In contrast, our method is fully end-to-end, eliminating the need for additional training processes and completely removing the reliance on feature extractor initialization.
II-C Self-Supervised Person Re-Identification
ReID [43, 44] aims at matching the same person image captured by non-overlapping cameras. Due to its critical role in video surveillance, ReID has attracted extensive research attention and achieved remarkable progress. However, these achievements are facilitated by extensive human-labeled data. To address this limitation, self-supervised learning methods [6, 7] have been proposed. Among them, clustering algorithms are widely employed to generate pseudo-labels, which serve as supervision signals for model training. In addition, benefiting from the introduction of instance-level contrastive learning [45] with memory banks and momentum update strategies [46], many methods [47, 48] consider each unlabeled sample as an individual class to learn discriminative instance-level feature representations. However, these instance-level features was modified largely during training, resulting in unstable representations [49]. To address this issue, recent studies [50, 51] have begun exploring cluster-level relationships to discover more stable and robust associations among pedestrian representations. Cluster Contrast [17] addresses the problem of cluster inconsistency by performing clustering and introducing a distinct representation for each cluster. Nevertheless, in cross-modal ReID scenarios, where images are captured by heterogeneous sensors, modality discrepancies hinder the effectiveness of these methods in learning modality-invariant features. ADCA[18] introduced a dual-path contrastive framework to bridge modality gaps and enhance cross-modal feature alignment. Building on this, SDCL [52] employs a dual-path Transformer for shallow-to-deep contrastive learning, and refines pseudo-labels via consistency constraints across feature hierarchies, leading to more stable and discriminative cross-modal representations.
The above methods focus on alleviating the dependency on human-labeled data in the ReID and provide valuable insights for UVGL. However, UVGL suffers from more severe viewpoint and scale variations. Furthermore, the spatial continuity of geographic environments inevitably induces feature ambiguity, severely degrading clustering effectiveness. Thereby, we incorporate instance-level and cluster-level features into a self-supervised UVGL method to learn more discriminative and robust cross-view representations.
III METHOD
As illustrated in Fig. 3, the proposed framework is composed of two key modules, DHML and ICEL. Additionally, a PLE strategy is employed to optimize the self-supervised learning process. Specifically, we adopt a weight-shared ConvNeXt-Tiny network as the backbone, equipped with cluster memory and a dual-path contrastive learning strategy to learn intra-view feature representations. To capture the dynamic variations of intra-class features caused by changes in viewpoint and scale, the DHML module establishes a collaborative relationship between short-term and long-term memory, thereby enhancing the discriminability of intra-view feature representations. Furthermore, the ICEL module leverages instance memory to construct neighborhood-driven dynamic constraints. By integrating neighborhood consistency with mutual information optimization, it guides cross-view features to form unified and consistent representations.
Problem Formulation: In the UVGL, we are given a set of paired UAV-satellite images, denoted by . Here, represents UAV-view images and represents satellite-view images. Unlike the supervised setting, our self-supervised method is developed without using paired UAV-satellite images, where the correspondences between UAV-view and satellite-view images are entirely unavailable. Thereby, we randomly sample UAV and satellite images independently during training. The objective is to discover reliable cross-view correspondences under such conditions and accomplish the UVGL task.
III-A Dual-Path Contrastive Learning Baseline
Cluster-based contrastive learning methods have achieved remarkable success in various fields. However, in UVGL, due to the unique challenges of cross-view scenarios, single-path cluster-based contrastive learning struggles to effectively capture view-specific features. To address this issue, we design the baseline network with a dual-path architecture, where UAV-view and satellite-view features are extracted using a shared backbone and clustered to assign pseudo-labels. Subsequently, a dual-path contrastive learning strategy [53] is employed to enhance intra-view representation learning.
We first introduce the notations used in this paper for better clarity. Let denote a set of drone images with instances, and denote a set of satellite images with instances. and represent the feature representations extracted by the shared feature extractor for the drone and satellite views, respectively. and denote the drone query instance features and satellite query instance features extracted by the feature extractor . The processing procedure of the backbone can be described as
Cross-View Feature Memory Initialization. At the beginning of each training epoch, the feature memories for both views are initialized. Specifically, the cluster representations of the drone view and the satellite view are stored in the drone memory dictionary and satellite memory dictionary , respectively, for view-specific memory initialization. This process can be formalized as
where , , , , denotes the -th cluster set in drone or satellite view, indicates the number of instances per cluster.
Cross-View Feature Memory Updating. During the training phase, we randomly sample distinct localization points from the training set, with each point comprising a fixed number of instances. This sampling strategy results in a minibatch containing a total of query images. Subsequently, the query features extracted from the minibatch are integrated into the cluster representations via a momentum-based update mechanism. Through this method, we iteratively update the features for both perspectives. This process can be formalized as
where is the momentum updating factor. is the iteration number.
Cross-View Contrastive Loss. Given drone and satellite query and , we compute the contrastive loss for drone view and satellite view by the following equations:
where denotes the temperature hyper-parameter. The cluster centroids are utilized as feature vectors at the cluster level to compute the distances between the query instance and all clusters. Here, corresponds to the positive cluster feature associated with .
The final optimization for baseline is denoted by the following combination:
III-B Dynamic Hierarchical Memory Learning
Geographic environments inherently exhibit spatial continuity, where neighboring regions usually present similar visual patterns, while distant regions display significant differences. However, such spatial characteristics often introduce ambiguity in intra-view feature representations. To address this, we propose a DHML module that jointly exploits long-term and short-term memory to capture spatial dynamics and enhance the discriminability of intra-view features.
Consistent with the baseline network, we construct the dynamic hierarchical memory learning in a dual-path manner, using drone feature extraction as an illustrative example. To build the long-term memory, we first construct two novel clustering memory dictionaries according to Eq.2, which are used to store the short-term and long-term memories, respectively.
where, represent the short-term memory embeddings that encode the cluster representations of the drone view, and represent the long-term memory embeddings that store historically accumulated stable representations of the drone view.
Unlike the weight updating scheme in Eq.4 and Eq.5, we assign different weights to the input features and the historical memory during the long-term memory update. Specifically, since the momentum parameter is set to 0.2 throughout this study, we enhance the influence of historical memory during the initialization of long-term memory updates. The update procedure can defined as
where, represents the drone query instance features. This obtain a new dictionary of long-term memory .
Additional, we further introduce an adaptive weight updating mechanism to update the long-term memory. In detail, we compute the Euclidean distance between the current input and its corresponding long-term memory, then take the average over all samples in the batch. This average distance is then mapped to the range via the sigmoid function, yielding an adaptive update coefficient :
where, denotes the batch size and represents the sigmoid function. The short-term memory is subsequently updated as follows:
Then, the long-term and short-term memories are fused in a weighted manner to yield the combined feature representation:
where and denote the respective fusion weights. As a result, we obtain a comprehensive memory dictionary that effectively integrates long-term global consistency with localized short-term adaptability.
Given the drone query , we compute the contrastive loss by measuring their similarity to the cluster representations maintained in the memory dictionary by the following equations:
where denotes the temperature hyper-parameter. The combined cluster centroids are utilized as feature vectors at the cluster level to compute the distances between the query instance and all clusters. Here, corresponds to the positive cluster feature associated with .
Finally, we introduce a fusion factor to perform weighted fusion of and , aiming to enhance the discriminability of intra-view feature representations and obtain the optimized loss . Similarly, can be obtained in the same way. Therefore, the dynamic hierarchical memory learning process can be summarized as
III-C Information Consistency Evolution Learning
The baseline network and the DHML module primarily focus on intra-view feature learning, lacking the ability to model and associate cross-view features effectively. Therefore, the proposed ICEL module employs a threshold-based and ranking-based neighborhood selection strategy to achieve effective cross-view feature alignment. By promoting the consistency evolution of cross-view feature representations, ICEL enhances both the discriminability and representation capability of cross-view features.
In the ICEL module, unlike the baseline and DHML module, we employ instance-level features instead of cluster-level representations. To achieve this, we define two instance memory dictionaries, and , which store the instance features extracted by the shared feature encoder, as described in Section III-A. This process can be described as
where and denote the drone and satellite instance features stored in memory dictionaries and , respectively.
In addition, we define the similarity between a given drone (or satellite) query and each drone (or satellite) instance in the training set as
where and are the indices of queries and instances. Thus, we obtain intra-view similarity , and cross-view similarity , .
Then, we adopt a threshold-based neighborhood selection strategy to select reliable neighbors, which can be defined as
where represents the set of drone feature instances selected as valid intra-view neighbors. denotes the set of drone feature instances that exhibit strong cross-view associations with satellite features. is a predefined threshold. The and can be obtained by the same way.
Additionally, we introduce a ranking-based neighborhood selection strategy. Specifically, we independently select the top- most relevant samples to construct a strictly filtered neighborhood, ensuring high-confidence associations, while also selecting the top- samples define an expanded neighborhood, capturing additional informative relationships. The process can be defined as
where , and , denote the selected intra-view and cross-view neighborhood sets, respectively, with and containing exactly elements, and and containing exactly elements. Similarly, the neighborhood sets , , , and can be obtained following the same selection strategy.
Based on the constructed neighborhoods, the normalized similarity distribution is computed within the larger neighborhood to promote the aggregation of similar samples in both intra-view and cross-view scenarios. This distribution is then optimized by a cross-entropy loss to enhance alignment between query samples and their neighbors. The intra-view and cross-view losses can be formulated as
where represents the batch size of the query , and denote the total number of neighborhood samples within the set and .
Incorporating neighborhoods , and , , we introduce consistency loss to enhance alignment of relative similarity distributions within the large-scale neighborhood across both intra-view and cross-view representations. The intra-view loss and cross-view loss under this consistency loss can be formulated as
where is the softmax-normalized similarity distribution, is used for summation to normalize the similarity scores over all neighborhood samples. and can be obtained in the same way.
Within a more restrictive and smaller set of neighbors, a mutual information constraint is proposed to learn intra-view and cross-view features representations. This complements the consistency constraint described above, achieving a balance between global structural alignment and local discriminative information. Within the selected neighborhood, we first compute, for each neighbor sample (where ), the similarity between the deep query feature and the neighbor feature using the similarity function . Based on these similarity scores, the posterior probability for each neighbor is defined via a softmax normalization as:
where denote the total number of neighborhood samples within the set and .
Therefore, the intra-view and cross-view mutual information losses and can be derived and are defined as follows:
where and represent the normalized similarity distributions over the selected intra-view and cross-view neighbors, respectively. and can be obtain in the same way.
For each view combination, the loss is computed as the weighted sum of three components: the original neighborhood loss , the strict neighborhood loss , and the extended neighborhood loss :
where and denote the drone or satellite view.
The total neighbor learning loss is then formulated as
Finally, the overall loss of our method is defined as
III-D Pseudo-Label Enhancement
In self-supervised learning, the quality of pseudo-labels exerts a decisive influence on model performance. Inspired by [52], our proposes a pseudo-label optimization strategy based on dynamic feature robustness. Specifically, we introduce small perturbations to the original feature representations and exploit the top-K neighbor consistency between the original and perturbed features to identify high-confidence samples while filtering out noisy instances. Then, a spatial smoothing operation based on the similarity matrix is applied to refine the pseudo-labels, enforcing local label consistency.
We first employe small Gaussian noise into the original satellite instances feature and the original drone instances features to obtain perturbed features, which are used for more robust similarity estimation and subsequent cluster consistency evaluation. It can be formulated as
where, with the similarities and , we can get two cross-view ranking lists and for and , which are denoted as and .
The label of the -th similar infrared instance in two ranking lists can be represented as
where and are refined -th cross-view labels of and through the ranking on cross-view initial features and perturbed features similarities.
So, we associate the cross-modality labels with the intersection of two label sets to investigate collaborative ranking consistency. This process can be described as
where records the samples with the same identity of top- identity in instances from top-1 to top-.
The reliable refined cross-view label of can be expressed by the label of maximum count in .
where denotes the counting function. The operation traverses and finds the labels with maximum number as cross-view refined labels.
Then, we convert the label list to the form of one-hot code matrix by setting the column according to the refined labels to 1 and the rest to 0.Then,two initial and perturbed homogeneous initial and deep similarity matrices are constructed to investigate intra-view ranking consistency, enhancing the precision of refined cross-view pseudo labels by:
where represents the intra-view similarity structure.
In order to explore the intra-view initial and perturbed collaboration, we calculate the sum of and by:
where indicates the consistency of initial and perturbed similarity matrix.
We keep the 5-max values of in each row to 1 and set the rest to 0, acquiring the ranking relations. The process of intra-modality ranking smoothing is formulated as follows:
where is the final refined cross-view label matrix of satellite instance. In , the column number of the maximum value in each row is the refined label of samples.
| University-1652 | |||||||
| DroneSatellite | SatelliteDrone | ||||||
| Model | Venue | Supervision Ratio | Paramars(M) | R@1 | AP | R@1 | AP |
| MuSe-Net[54] | PR’2024 | 100(%) | 50.47 | 74.48 | 77.83 | 88.02 | 75.10 |
| MCCG[1] | TCSVT’2023 | 100(%) | 56.65 | 89.40 | 91.07 | 95.01 | 89.93 |
| SDPL[3] | TCSVT’2024 | 100(%) | 42.56 | 90.16 | 91.64 | 93.58 | 89.45 |
| CCR[32] | TCSVT’2024 | 100(%) | 156.57 | 92.54 | 93.78 | 95.15 | 91.80 |
| Sample4Geo[29] | ICCV’2023 | 100(%) | 87.57 | 92.65 | 93.81 | 95.14 | 91.39 |
| SRLN[31] | TGRS’2024 | 100(%) | 193.03 | 92.70 | 93.77 | 95.14 | 91.97 |
| MEAN[2] | arXiv’2024 | 100(%) | 36.50 | 93.55 | 94.53 | 96.01 | 92.08 |
| DAC[4] | TCSVT’2024 | 100(%) | 96.50 | 94.67 | 95.50 | 96.43 | 93.79 |
| Ours | - | 100(%) | 28.59 | 92.02 | 93.36 | 95.57 | 91.27 |
| EM-CVGL[5] | TGRS’2024 | 0(%) | 2.25 | 70.29 | 74.93 | 79.03 | 61.03 |
| CDIKTNet[42] | arXiv’2025 | 0(%) | 91.52 | 83.30 | 85.73 | 87.73 | 76.53 |
| Ours | - | 0(%) | 28.59 | 90.17 | 91.67 | 95.01 | 88.95 |
IV Experimental results
IV-A Experimental Datasets and Evaluation Metrics
To evaluate our self-supervised method, we conduct experiments on three benchmark datasets including University-1652 [12], SUES-200 [33] and DenseUAV [54].
University-1652 contains images from 1,652 buildings across 72 universities worldwide. It includes drone, satellite, and ground-level views. The training set consists of 701 buildings from 33 universities, and the test set covers 951 buildings from 39 geographically disjoint universities.
SUES-200 focuses on altitude diversity in aerial imagery, containing 200 locations with 120 for training and 80 for testing. Each location provides one satellite image and aerial images captured at four altitudes (150m, 200m, 250m, and 300m), covering urban areas, natural landscapes, and water bodies. SUES-200 evaluates cross-view retrieval under varying flight heights, reflecting real-world UAV scenarios.
DenseUAV focuses on UAV self-positioning in low-altitude urban environments, containing over 27,000 UAV and satellite images collected from 14 university campuses. It features dense sampling, multi-scale satellite images, and varying altitudes (80m, 90m, and 100m) under multiple temporal conditions. The training set includes 6,768 UAV images and 13,536 satellite images, while the test set comprises 2,331 UAV images and 4,662 satellite images from non-overlapping areas. DenseUAV poses challenges such as cross-view matching, neighbor confusion, and spatial-temporal variation.
We evaluate the models using Recall@K (R@K) and Average Precision (AP), where R@K measures the proportion of correct matches among the Top-results and AP balances precision and recall. Model efficiency is assessed by the number of parameters, reflecting portability under resource constraints. All comparisons are conducted using the best configurations for each method.
IV-B Implementation Details
A ConvNeXt-Tiny model, pre-trained on ImageNet, serves as the backbone for feature extraction, with a newly added classifier module initialized via the Kaiming method. During both training and testing, all input images are resized to . We employ several data augmentation techniques, including random cropping, random horizontal flipping, and random rotation. The batch size is set to 64, and DBSCAN [8] is used to generate pseudo labels. For optimization, we use the SGD optimizer with an initial learning rate of 0.001, training the model for 30 epochs in total. All experiments are conducted under the PyTorch framework on an Ubuntu 22.04 system equipped with four NVIDIA RTX 4090 GPUs. Additionally, in the University-1652 and SUES-200 datasets,we replicate each satellite feature 50 times for clustering purposes, leaving UAV features unchanged. In DenseUAV, we replicate UAV features four times at each altitude, and replicate the satellite features three times for each time point and altitude.
IV-C Comparison with State-of-the-art Methods
To evaluate the effectiveness of our method, we compare it with the state-of-the-art supervised methods MuSe-Net [54], MCCG [1], SDPL [3], CCR [32], Sample4Geo [29], SRLN[31], MEAN [2] and DAC [4], and the recent self-supervised methods EM-CVGL [5], CDIKTNet [42].
Results on University-1652. As shown in Table I, our method exhibits superior performance compared with the current advanced self-supervised methods. Specifically, our method achieves R@1/AP scores of 90.17%/91.67% and 95.01%/88.95% under the DroneSatellite and SatelliteDrone settings, respectively. Compared with EM-CVGL, our method achieves significant improvements, outperforming it by approximately 20% in R@1 and 17% in AP on the DroneSatellite setting, and by 16% in R@1 and 28% in AP on the SatelliteDrone setting. Furthermore, even compared with CDIKTNet, which utilizes paired data from other domains for self-supervised fine-tuning, our method still consistently achieves superior performance without relying on any cross-domain paired data.
We further benchmark against eight representative supervised methods, and the results show that our method clearly surpasses some of them, including MuSe-Net, MCCG, and SDPL. Furthermore, with self-supervised pre-training and subsequent supervised fine-tuning, our method achieves competitive performance under the supervised setting, comparable to CCR, Sample4Geo, and SRLN.
| SUES-200 | |||||||||||
| DroneSatellite | |||||||||||
| 150m | 200m | 250m | 300m | ||||||||
| Model | Venue | Supervision Ratio | Paramars(M) | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
| MCCG[1] | TCSVT’2023 | 100(%) | 56.65 | 82.22 | 85.47 | 89.38 | 91.41 | 93.82 | 95.04 | 95.07 | 96.20 |
| SDPL[3] | TCSVT’2024 | 100(%) | 42.56 | 82.95 | 85.82 | 92.73 | 94.07 | 96.05 | 96.69 | 97.83 | 98.05 |
| CCR[32] | TCSVT’2024 | 100(%) | 156.57 | 87.08 | 89.55 | 93.57 | 94.90 | 95.42 | 96.28 | 96.82 | 97.39 |
| SRLN[31] | TGRS’2024 | 100(%) | 193.03 | 89.90 | 91.90 | 94.32 | 95.65 | 95.92 | 96.79 | 96.37 | 97.21 |
| Sample4Geo[29] | ICCV’2023 | 100(%) | 87.57 | 92.60 | 94.00 | 97.38 | 97.81 | 98.28 | 98.64 | 99.18 | 99.36 |
| DAC[4] | TCSVT’2024 | 100(%) | 96.50 | 96.80 | 97.54 | 97.48 | 97.97 | 98.20 | 98.62 | 97.58 | 98.14 |
| MEAN[2] | arXiv’2024 | 100(%) | 36.50 | 95.50 | 96.46 | 98.38 | 98.72 | 98.95 | 99.17 | 99.52 | 99.63 |
| Ours | - | 100(%) | 28.59 | 94.00 | 95.21 | 97.83 | 98.88 | 99.14 | 95.44 | 99.25 | 99.39 |
| EM-CVGL[5] | TGRS’2024 | 0(%) | 2.25 | 55.23 | 60.80 | 60.95 | 61.03 | 68.10 | 72.62 | 74.42 | 78.20 |
| CDIKTNet[42] | arXiv’2025 | 0(%) | 91.52 | 82.75 | 85.25 | 89.35 | 91.13 | 93.15 | 94.39 | 95.18 | 96.12 |
| Ours† | - | 0(%) | 28.59 | 82.60 | 85.60 | 90.48 | 92.28 | 94.33 | 95.44 | 96.78 | 97.42 |
| Ours‡ | - | 0(%) | 28.59 | 87.00 | 89.20 | 93.98 | 95.11 | 96.50 | 97.23 | 97.23 | 97.86 |
| Ours∗ | - | 0(%) | 28.59 | 90.98 | 92.66 | 95.03 | 96.14 | 97.45 | 98.03 | 98.23 | 98.64 |
| SatelliteDrone | |||||||||||
| 150m | 200m | 250m | 300m | ||||||||
| Model | Venue | Supervision Ratio | Paramars(M) | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
| CCR[32] | TCSVT’2024 | 100(%) | 156.57 | 92.50 | 88.54 | 97.50 | 95.22 | 97.50 | 97.10 | 97.50 | 97.49 |
| Sample4Geo[29] | ICCV’2023 | 100(%) | 87.57 | 92.60 | 94.00 | 97.38 | 97.81 | 98.28 | 98.64 | 99.18 | 99.36 |
| MCCG[1] | TCSVT’2023 | 100(%) | 56.65 | 93.75 | 89.72 | 93.75 | 92.21 | 96.25 | 96.14 | 98.75 | 96.64 |
| SDPL[3] | TCSVT’2024 | 100(%) | 42.56 | 93.75 | 83.75 | 96.25 | 92.42 | 97.50 | 95.65 | 96.25 | 96.17 |
| SRLN[31] | TGRS’2024 | 100(%) | 193.03 | 93.75 | 93.01 | 97.50 | 95.08 | 97.50 | 96.52 | 97.50 | 96.71 |
| DAC[4] | TCSVT’2024 | 100(%) | 96.50 | 97.50 | 94.06 | 98.75 | 96.66 | 98.75 | 98.09 | 98.75 | 97.87 |
| MEAN[2] | arXiv’2024 | 100(%) | 36.50 | 97.50 | 94.75 | 100.00 | 97.09 | 100.00 | 98.28 | 100.00 | 99.21 |
| Ours | - | 100(%) | 28.59 | 97.50 | 92.71 | 97.50 | 97.43 | 100.00 | 98.98 | 98.75 | 98.25 |
| EM-CVGL[5] | TGRS’2024 | 0(%) | 2.25 | 73.75 | 54.00 | 91.25 | 65.65 | 96.25 | 72.02 | 97.50 | 74.74 |
| CDIKTNet[42] | arXiv’2025 | 0(%) | 91.52 | 88.75 | 80.33 | 93.75 | 88.17 | 95.00 | 92.15 | 98.75 | 94.37 |
| Ours† | - | 0(%) | 28.59 | 91.25 | 81.28 | 97.50 | 90.62 | 98.75 | 94.74 | 97.50 | 96.17 |
| Ours‡ | - | 0(%) | 28.59 | 95.00 | 85.46 | 97.50 | 92.72 | 100.00 | 95.43 | 100.00 | 96.39 |
| Ours∗ | - | 0(%) | 28.59 | 96.25 | 90.05 | 97.50 | 94.80 | 98.75 | 96.82 | 98.75 | 97.33 |
Results on SUES-200. As shown in Table II, As shown in Table II, our method, when directly transferring the pre-trained model from the University-1652 dataset to the SUES-200 dataset without any re-training, consistently outperforms all existing self-supervised methods and achieves competitive performance compared to supervised methods such as MCCG and SDPL. Moreover, by performing self-supervised re-training on the SUES-200 dataset, our method brings a significant performance boost, achieving results comparable to the supervised CCR. Notably, when further fine-tuning the transferred model for only one epoch, we observe additional improvements, surpassing the supervised SRLN. These results clearly demonstrate the effectiveness of our method in self-supervised learning, transfer learning, and efficient fine-tuning.
Results on Cross-Domain Generalization Performance: Our method consistently outperforms existing self-supervised method under cross-domain settings, showing both significant improvements and stable generalization. In the Drone to Satellite setting, our model achieves better performance than EM-CVGL at all altitudes. At 150 meters, R@1 and AP increase by approximately 22% and 20%; at 200 meters, the gains are around 19% and 16%; at 250 meters, nearly 14% and 12%; and even at 300 meters, our method still outperforms by about 11% and 9%. In the Satellite to Drone scenario, similar improvements are observed. At 150 meters, R@1 and AP improve by approximately 18% and 25%; at 200 meters, by 10% and 20%; at 250 meters, by 6% and 13%; and at 300 meters, by 3% and 10%. These results confirm the robustness and strong cross-domain generalization capability of our method across varying altitudes.Remarkably, our method even surpasses all existing state-of-the-art supervised methods in terms of cross-domain generalization. These results confirm the robustness and strong cross-domain generalization capability of our method across varying altitudes.
| University-1652SUES-200 | |||||||||||
| DroneSatellite | |||||||||||
| 150m | 200m | 250m | 300m | ||||||||
| Model | Venue | Supervision Ratio | Paramars(M) | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
| MCCG[1] | TCSVT’2023 | 100(%) | 56.65 | 57.62 | 62.80 | 66.83 | 71.60 | 74.25 | 78.35 | 82.55 | 85.27 |
| Sample4Geo[29] | ICCV’2023 | 100(%) | 87.57 | 70.05 | 74.93 | 80.68 | 83.90 | 87.35 | 89.72 | 90.03 | 91.91 |
| DAC[4] | TCSVT’2024 | 100(%) | 96.50 | 76.65 | 80.56 | 86.45 | 89.00 | 92.95 | 94.18 | 94.53 | 95.45 |
| CAMP[30] | TGRS’2024 | 100(%) | 91.40 | 78.90 | 82.38 | 86.83 | 89.28 | 91.95 | 93.63 | 95.68 | 96.65 |
| MEAN[2] | arXiv’2024 | 100(%) | 36.50 | 81.73 | 87.72 | 89.05 | 91.00 | 92.13 | 93.60 | 94.63 | 95.76 |
| EM-CVGL[5] | TGRS’2024 | 0(%) | 2.25 | 60.03 | 65.69 | 71.50 | 76.17 | 80.35 | 83.85 | 85.93 | 88.34 |
| Ours | - | 0(%) | 28.59 | 82.60 | 85.60 | 90.48 | 92.28 | 94.33 | 95.44 | 96.78 | 97.42 |
| SatelliteDrone | |||||||||||
| 150m | 200m | 250m | 300m | ||||||||
| Model | Venue | Supervision Ratio | Paramars(M) | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
| MCCG[1] | TCSVT’2023 | 100(%) | 56.65 | 61.25 | 53.51 | 82.50 | 67.06 | 81.25 | 74.99 | 87.50 | 80.20 |
| Sample4Geo[29] | ICCV’2023 | 100(%) | 87.57 | 83.75 | 73.83 | 91.25 | 83.42 | 93.75 | 89.07 | 93.75 | 90.66 |
| DAC[4] | TCSVT’2024 | 100(%) | 96.50 | 87.50 | 79.87 | 96.25 | 88.98 | 95.00 | 92.81 | 96.25 | 94.00 |
| CAMP[30] | TGRS’2024 | 100(%) | 91.40 | 87.50 | 78.98 | 95.00 | 87.05 | 95.00 | 91.05 | 96.25 | 93.44 |
| MEAN[2] | arXiv’2024 | 100(%) | 36.50 | 91.25 | 81.50 | 96.25 | 89.55 | 95.00 | 92.36 | 96.25 | 94.32 |
| EM-CVGL[5] | TGRS’2024 | 0(%) | 2.25 | 73.75 | 56.99 | 87.50 | 70.62 | 92.50 | 81.18 | 95.00 | 86.04 |
| Ours | - | 0(%) | 28.59 | 91.25 | 81.28 | 97.50 | 90.62 | 98.75 | 94.74 | 97.50 | 96.17 |
| DenseUAV | |||||||||||
| All height | 80m | 90m | 100m | ||||||||
| Model | Venue | Supervision Ratio | Paramars(M) | R@1 | R@5 | R@1 | R@5 | R@1 | R@5 | R@1 | R@5 |
| ConvNext-Tiny[34] | TIP’2024 | 100(%) | 30.10 | 60.23 | 81.94 | - | - | - | - | - | - |
| Sample4Geo[29] | ICCV’2023 | 100(%) | 87.57 | 80.57 | 96.53 | 80.57 | 96.53 | 86.87 | 98.71 | 91.38 | 99.74 |
| CAMP[30] | TGRS’2024 | 100(%) | 91.40 | 88.72 | 97.76 | 76.45 | 96.01 | 84.81 | 99.23 | 89.19 | 99.23 |
| DAC[4] | TCSVT’2024 | 100(%) | 96.50 | 84.47 | 96.53 | 80.92 | 97.04 | 85.32 | 98.71 | 85.46 | 99.10 |
| MEAN[2] | arXiv’2024 | 100(%) | 36.50 | 90.18 | 97.86 | 78.76 | 97.04 | 85.33 | 99.10 | 89.32 | 98.97 |
| Ours | - | 100(%) | 28.59 | 86.31 | 97.81 | 86.10 | 98.46 | 88.93 | 99.61 | 90.35 | 99.49 |
| EM-CVGL[5] | TGRS’2024 | 0(%) | 2.25 | 18.15 | 50.97 | 18.02 | 50.36 | 17.76 | 50.19 | 17.50 | 49.98 |
| CDIKTNet[42] | arXiv’2025 | 0(%) | 91.52 | 22.57 | 40.24 | 17.25 | 40.15 | 23.04 | 48.39 | 25.48 | 52.38 |
| Ours | - | 0(%) | 28.59 | 73.79 | 90.69 | 73.10 | 92.92 | 74.65 | 95.37 | 75.03 | 95.50 |
Results on DenseUAV: The DenseUAV dataset presents a highly challenging scenario due to its densely sampled nature, where neighboring images exhibit high visual similarity, often leading to severe confusion in UVGL. In addition, the dataset includes satellite images captured at different times and UAV images from multiple altitudes, which further increases intra-domain variability and task complexity. As shown in Table IV, under the fully unsupervised setting, our method achieves 73.79% R@1 and 90.69% R@5 across all altitudes, significantly outperforming existing self-supervised methods. It is noteworthy that the two representative self-supervised baselines consistently yield poor performance on DenseUAV, highlighting their limitations in handling densely sampled and spatiotemporal diverse scenarios. Furthermore, after supervised fine-tuning, our method surpasses several supervised baselines, including ConvNeXt-Tiny, Sample4Geo, CAMP, and DAC. In some cases, it even outperforms the current best-performing supervised method, MEAN. These results clearly demonstrate the effectiveness of our method in addressing extreme feature confusion caused by spatial continuity and appearance shifts induced by temporal variation, confirming its strong generalization and robustness.
IV-D Ablation Studies
To thoroughly evaluate the effectiveness of our proposed method, we conduct an ablation study on the University-1652 dataset. We systematically assess the impact of each individual module and investigate their contributions under different configurations. The results are presented in Table LABEL:Ablation_Studies.
ConvNeXt-Tiny Backbone. The Backbone refers to the untrained ConvNeXt-Tiny network used as a direct feature extractor. As shown in Table LABEL:Ablation_Studies, its performance is significantly limited, achieving an R@1 of only 9.55% and 35.24%, and an AP of 11.88% and 13.20% for the Drone Satellite and Satellite Drone, respectively. These results indicate that the backbone alone lacks the ability to extract discriminative cross-view features.
Effectiveness of Dual-Path Contrastive Learning Baseline. After incorporating the dual-path contrastive learning baseline, the R@1 improves to 71.06% and 81.31%, while AP reaches 74.91% and 65.26%, respectively. Although the baseline primarily learns intra-view feature discrimination, the training process also facilitates a certain degree of alignment between UAV and satellite representations. This leads to the formation of a more consistent and unified feature representation space, which serves as a solid foundation for subsequent cross-view learning.
Effectiveness of Dynamic Hierarchical Memory Learning. Compared with the contrastive learning baseline, incorporating DHML leads to significant performance improvements under both settings. Specifically, R@1 increases by 12.97% and 9.13%, and AP improves by 11.5% and 17.48% , reaching 84.03% and 90.44% in R@1, and 86.41% and 82.74% in AP, respectively. These gains can be attributed to the joint modeling of long-term and short-term feature dependencies introduced by DHML, which enhances the capacity to capture spatial structures in UVGL scenarios and enhances intra-view feature representation and discrimination.
| Setting | University-1652 | |||||||
| DroneSatellite | SatelliteDrone | |||||||
| Backbone | Baseline | DHML | ICEL | PLE | R@1 | AP | R@1 | AP |
| ✓ | 9.55 | 11.88 | 35.24 | 13.20 | ||||
| ✓ | ✓ | 71.06 | 74.91 | 81.31 | 65.26 | |||
| ✓ | ✓ | ✓ | 84.03 | 86.41 | 90.44 | 82.74 | ||
| ✓ | ✓ | ✓ | 74.48 | 77.96 | 89.87 | 74.37 | ||
| ✓ | ✓ | ✓ | ✓ | 86.13 | 88.31 | 94.01 | 82.97 | |
| ✓ | ✓ | ✓ | ✓ | ✓ | 90.17 | 91.67 | 95.01 | 88.95 |
Effectiveness of Information Consistency Evolution Learning. Although DHML enhances intra-view feature representation and discrimination, its optimization remains limited to within-view information. To address this limitation, ICEL is proposed to learn cross-view dependencies. By enabling bidirectional information interaction between different views, ICEL facilitates cross-view feature refinement. As a result, R@1 improves to 86.13% and 51.04%, while AP reaches 88.31% and 82.97%. These results further validate the effectiveness of ICEL in enhancing cross-view consistency for UVGL, thereby contributing to overall performance improvement.
Pseudo-Label Enhancement (PLE): Refining Feature Learning with Robust Labels. To mitigate pseudo-label noise caused by clustering, we apply PLE to improve label reliability. After incorporating PLE, our final model achieves state-of-the-art performance, with R@1 increasing to 90.17% and 95.01%, and AP reaching 91.67% and 88.95%. This confirms that refining pseudo-label quality significantly enhances the final model’s effectiveness in cross-geographic view matching.
The ablation study systematically evaluates the contribution of each component in the proposed self-supervised framework. Starting with a backbone that lacks sufficient discriminative power, we first establish a dual-branch contrastive learning baseline. On top of this, DHML is introduced to strengthen intra-view feature representation and discriminability, ICEL is employed to facilitate cross-view feature association, and PLE is leveraged for pseudo-label refinement. Experimental results show that our method significantly enhances the learning of discriminative and view-consistent representations under self-supervised learning.
IV-E Visualization
Feature space and similarity distribution. To further illustrate the effectiveness of the our method, We perform feature space (t-SNE[55] map) and similarity distribution visualization for our method, as presented in Fig 4.Comparing (a) Baseline and (b) Ours, it is evident that our method enhances the clustering of cross-view positive samples, effectively bringing the drone and satellite representations of the same location closer together. Additionally, the separation between different locations is more distinct, indicating better feature discriminability. The second row shows the similarity distribution between drone-satellite positive and negative pairs. Our method demonstrates a clearer distinction between positive and negative pairs, with a reduced overlap between distributions, which suggests improved cross-view matching and reduced feature ambiguity. This highlights the effectiveness of our method in UVGL.
Cluster analysis. To validate the efficacy of our method, as shown in Fig. 5, we illustrate the evolution of cluster counts across epochs for three cross-view datasets. In the DenseUAV dataset (2,255 locations), satellite images exhibit significant clustering fluctuations during the early training stages, due to low intra-class discriminability and weak cross-view correlations. In contrast, the University-1652 dataset (701 locations) demonstrates greater UAV image clustering instability, while satellite image clusters stabilize more rapidly in the initial epochs. The SUES-200 dataset (1,200 locations), with its smaller data size and simpler distribution, exhibits minimal early-stage variation and stabilizes faster than larger datasets. As training progresses, the cluster counts across all datasets gradually converge toward the ground-truth location counts, despite initial inter-class confusion. This process validates the ability of our method to progressively learn discriminative representations.
Retrieval results. To demonstrate the effectiveness of our method, we present retrieval results on the University-1652 and DenseUAV dataset in Fig. 6 and Fig. 7. Green borders indicate correct matches, while red borders denote errors, showing that our method significantly improves retrieval accuracy.
In the University-1652 dataset, the dronesatellite retrieval task presents drone-view images as queries and retrieves corresponding satellite-view images. Since each location is associated with a single satellite image, our method consistently achieves Top-1 accuracy, demonstrating its high effectiveness in cross-view matching. In the satellitedrone setting, our method accurately retrieves the corresponding drone-view images given satellite queries, further showcasing its strong cross-view localization capability and reliable feature alignment across perspectives.
In the DenseUAV dataset, the retrieval task becomes significantly more challenging due to dense spatial sampling, multi-scale satellite imagery, and temporal dynamics, all of which exacerbate the difficulty of cross-view representation learning. In the Drone Satellite retrieval setting, despite variations in UAV altitudes and satellite image resolutions, our method effectively captures spatial-geographic cues, achieving high retrieval accuracy while substantially reducing mismatches. Compared to University-1652, DenseUAV provides a more realistic and complex benchmark for UAV-based geo-localization. Under this setting, our method exhibits strong cross-view retrieval performance, demonstrating robust feature discrimination in densely distributed urban regions.
V CONCLUSION
This paper investigates a highly valuable and challenging task namely self-supervised UAV-view cross-view geo-localization, aiming to reduce the reliance on paired UAV–satellite images. To this end, we propose a self-supervised UVGL method, which employs a shallow backbone with only 9 layers for feature extraction. Pseudo labels are generated via clustering, and a dual-branch contrastive learning mechanism is introduced to enhance the capability in representing intra-view visual patterns. To further improve intra-view discriminability and consistency, we design a dynamic hierarchical memory learning module. Additionally, to mitigate the structural and semantic domain gap between views, we propose an information consistency evolution learning module, which facilitates semantic alignment by learning high-dimensional feature consistency across views. To improve the quality and robustness of pseudo labels, we introduce a bidirectional feature consistency-based Pseudo Label Enhancement strategy, which refines pseudo-supervision signals and boosts cross-view representation learning. As a result, the proposed method effectively unifies both intra-view and cross-view feature representations in the embedding space. Extensive experiments conducted on three public benchmarks University-1652, SUES-200, and DenseUAV, demonstrate that our method significantly outperforms existing self-supervised methods, even surpassing several fully supervised methods in certain scenarios, without relying on paired UAV–satellite images. These results validate the effectiveness and generalization capability of our method, highlighting its potential for real-world UVGL applications.

### v3

@@
-Without Paired Labeled Data: An End-to-End Self-Supervised Paradigm for UAV-View Geo-Localization
+Without Paired Labeled Data: End-to-End Self-Supervised Learning for Drone-view Geo-Localization
 Abstract
-UAV-View Geo-Localization (UVGL) aims to achieve accurate localization of unmanned aerial vehicles (UAVs) by retrieving the most relevant GPS-tagged satellite images.
-However, existing methods heavily rely on pre-paired UAV-satellite images for supervised learning.
-Such dependency not only incurs high annotation costs but also severely limits scalability and practical deployment in open-world UVGL scenarios.
-To address these limitations, we propose an end-to-end self-supervised UVGL method.
-Our method leverages a shallow backbone network to extract initial features, employs clustering to generate pseudo labels, and adopts a dual-path contrastive learning architecture to learn discriminative intra-view representations.
-Furthermore, our method incorporates two core modules, the dynamic hierarchical memory learning module and the information consistency evolution learning module.
-The dynamic hierarchical memory learning module combines short-term and long-term memory to enhance intra-view feature consistency and discriminability.
-Meanwhile, the information consistency evolution learning module leverages a neighborhood-driven dynamic constraint mechanism to systematically capture implicit cross-view semantic correlations, thereby improving cross-view feature alignment.
-To further stabilize and strengthen the self-supervised training process, a pseudo-label enhancement strategy is introduced, which refines the quality of pseudo supervision.
-Our method ultimately constructs a unified cross-view feature representation space under self-supervised settings.
+Drone-view Geo-Localization (DVGL) aims to achieve accurate localization of drones by retrieving the most relevant GPS-tagged satellite images.
+However, most existing methods heavily rely on strictly pre-paired drone-satellite images for supervised learning.
+When the target region shifts, new paired samples are typically required to adapt to the distribution changes.
+The high cost of annotation and the limited transferability of these methods significantly hinder the practical deployment of DVGL in open-world scenarios.
+To address these limitations, we propose an end-to-end self-supervised learning method with a shallow backbone network.
+It employs a clustering algorithm to generate pseudo-labels and adopts a dual-path contrastive learning framework to learn discriminative intra-view representations.
+Furthermore, our method incorporates two core modules, including the dynamic hierarchical memory learning module (DHML) and the information consistency evolution learning module (ICEL).
+The DHML combines short-term and long-term memory to enhance intra-view feature consistency and discriminability.
+Meanwhile, the ICEL module utilizes a neighborhood-driven dynamic constraint mechanism to systematically capture implicit cross-view semantic correlations, consequently improving cross-view feature alignment.
+To further stabilize and strengthen the self-supervised training process, a pseudo-label enhancement strategy is introduced to enhance the quality of pseudo supervision.
 Extensive experiments on three public benchmark datasets demonstrate that the proposed method consistently outperforms existing self-supervised methods and even surpasses several state-of-the-art supervised methods.
-These results validate the effectiveness and generalizability of our method, highlighting its potential for real-world deployment in UVGL.
 Our code is available at https://github.com/ISChenawei/DMNIL.
 Index Terms:
-UAV-View Geo-Localization, Self-Supervised Learning, Contrastive Learning.I Introduction
-UAV-View Geo-Localization (UVGL) aims to achieve high-precision localization of UAV in GPS-denied scenarios by retrieving GPS-tagged satellite images that correspond to UAV-captured near-ground view images [1].
+Drone-view Geo-Localization, Self-Supervised Learning, Contrastive Learning.I Introduction
+Drone-View Geo-Localization (DVGL) is intended to achieve high-precision localization of drones in GPS-denied scenarios by retrieving GPS-tagged satellite images that correspond to drone aerial images [1].
 This capability is critical for applications such as autonomous navigation, disaster rescue, and emergency response.
-In recent years, existing UVGL methods have primarily relied on supervised learning using strictly paired UAV-satellite images [2, 3, 4].
-Although these supervised methods have achieved encouraging success, they also reveal several inherent limitations: (1) acquiring instance-level paired UAV-satellite images is highly costly [5], leading to increased training overhead; (2) large amounts of unpaired cross-view image resources remain underutilized; and (3) when applied to new scenarios, these methods often require recollection and re-pairing UAV-satellite images for training to adapt to domain shifts.
-These limitations significantly hinder the scalability and practical deployment of UVGL in real-world scenarios.
-Thereby, it is necessary and of great significance to explore a self-supervised UVGL method that can effectively mine latent cross-view correlations from large-scale unpaired UAV-satellite images.
-In recent years, self-supervised person re-identification (ReID) methods [6, 7] have achieved remarkable progress, offering crucial theoretical foundations for UVGL.
+In recent years, existing DVGL methods have mainly relied on supervised learning using strictly paired drone-satellite images [2, 3].
+Despite the promising performance demonstrated by these supervised methods, several intrinsic limitations are exposed.
+Firstly, the acquisition of instance-level paired drone-satellite images incurs additional training expenses [4].
+Secondly, large amounts of unpaired cross-view image resources remain under-exploited.
+Lastly, when these methods are applied to new scenarios, the need arises for re-collection and re-pairing of drone-satellite images for training purposes to accommodate domain shifts.
+These limitations significantly hinder the practical deployment of DVGL in open-world scenarios.
+Thereby, it is necessary to explore a self-supervised method that can extract latent cross-view correlations from large-scale unpaired drone-satellite images.
+In recent years, self-supervised person re-identification (ReID) methods [5, 6] have achieved remarkable progress to offer crucial theoretical foundations for DVGL.
 Typically, these methods employ clustering algorithms [7, 8] on features extracted by a backbone network to generate pseudo-labels, which serve as the initial supervision signals for self-supervised learning.
-However, the quality of these pseudo-labels is heavily dependent on the discriminative capability of backbone network.
-As illustrated in Fig.
-1, compared with conventional ReID, UVGL faces additional challenges.
-These factors make it more difficult for the backbone to extract discriminative features, as illustrated in Fig.
-2.
-When pseudo-labels contain a high error rate and the model lacks the capacity to capture intra-view and cross-view consistency and discriminative features, the self-supervised learning process becomes prone to catastrophic negative optimization.
-In this paper, we propose a novel end-to-end self-supervised UVGL method.
-A lightweight, shallow ConvNeXt-Tiny backbone network [10] is employed to extract features, which are subsequently clustered to generate pseudo-labels.
-Combined with a dual-path contrastive learning strategy [17, 18], these components form the baseline network of our self-supervised method for UVGL to learn intra-view feature representations.
-Building upon this baseline, we proposed the dynamic hierarchical memory learning (DHML) module and the information consistency evolution learning (ICEL) module as two key components of our method.
-The DHML module captures the relationship between short-term and long-term memory to learn the discriminability of intra-view feature representations.
-The ICEL module constructs a neighborhood-driven dynamic constraint mechanism based on feature distributions.
-By integrating neighborhood consistency constraints with mutual information optimization, it effectively aligns cross-view feature representations within a shared embedding space, promoting both consistency and discriminative learning.
-Additionally, to improve the reliability of pseudo-labels, we introduce a pseudo-label enhancement (PLE) strategy, which contributes to a more stable and effective self-supervised feature learning process.
-The main contributions can be summarized as follows:
-We propose an end-to-end self-supervised method for UVGL.
-Built upon a shallow backbone and a dual-path contrastive learning strategy, our method integrates dynamic hierarchical memory learning and information consistency evolution learning to effectively capture and enhance feature consistency and discriminability for both intra-view and cross-view representation learning.
-We propose a dynamic memory hierarchical learning module that captures the relationships among different levels of memory to adaptively model intra-class and inter-class feature dynamics within the intra-view, thereby enhancing the discriminability of feature representations.
-we propose a novel information consistency evolution learning module which combines neighborhood consistency and mutual information optimization to learn consistent and discriminative cross-view representations.
-Extensive experiments demonstrate that our method, utilizing only a lightweight backbone, achieves state-of-the-art performance on three UVGL benchmarks, outperforming even certain supervised methods.
+However, as illustrated in Fig.
+1, compared to conventional ReID tasks, DVGL presents more severe challenges, such as large intra-view ambiguities and cross-view discrepancies.
+As shown in Fig.
+2, these issues hinder the backbone network from extracting stable and discriminative features at the early stage, which directly undermines the quality of pseudo-labels generated during the initial clustering phase.
+Without effective strategies to guide the learning of intra-view and cross-view consistency as well as feature discrimination, the model lacks the necessary ability to extract meaningful representations from low-quality pseudo-labels.
+As a result, the learning process may precipitate a detrimental feedback loop, wherein inaccurate pseudo-labels undermine the quality of feature representations to further compromise pseudo-label quality in subsequent iterations, ultimately leading to the collapse of self-supervised training.
+Although a current self-supervised method [9] proposed a coarse-to-fine framework for DVGL with a similar dual-path contrastive learning framework, it relies on explicit appearance-level transformations rather than intrinsically modeling the underlying cross-view feature relationships.
+In the shared space, it still exhibits limited cross-view alignment and susceptibility to feature noise.
+In this paper, we propose a novel end-to-end self-supervised learning method for DVGL.
+A lightweight and shallow ConvNeXt-Tiny backbone network [10] is employed to extract features, which are subsequently clustered to generate pseudo-labels.
+Combined with a dual-path contrastive learning strategy [17, 18], the backbone network coupled with a downstream clustering process forms the baseline network of our method for learning intra-view feature representations.
+Building upon this baseline, we design the dynamic hierarchical memory learning (DHML) module and the information consistency evolution learning (ICEL) module as two key components of our method.
+The DHML module captures the relationship between short-term and long-term memory to enhance the discriminability of intra-view feature representations.
+The ICEL module constructs a neighborhood-driven dynamic constraint mechanism by integrating neighborhood consistency constraints with mutual information optimization to facilitate cross-view feature alignment and discriminative representation learning.
+DHML enhances intra-view discrimination and ICEL improves cross-view consistency.
+Through joint optimization of intra-view and cross-view features, they collectively strengthen the feature discriminability and consistency.
+Additionally, a pseudo-label enhancement (PLE) strategy is employed to improve the reliability of pseudo-labels and guide more stable self-supervised feature learning.
+The main contributions can be summarized as follows.
+We propose an end-to-end self-supervised learning method for DVGL.
+Built upon a shallow backbone and a dual-path contrastive learning strategy, our method integrates dynamic hierarchical memory learning and information consistency evolution learning to enhance feature consistency and discriminability for both intra-view and cross-view representation learning.
+Additionally, a pseudo-label enhancement (PLE) strategy is employed to guide more stable self-supervised learning.
+To address the intra-view ambiguity caused by spatial continuity across geographic viewpoints, we propose a dynamic memory hierarchical learning (DHML) module that captures the dependencies between different memory levels to adaptively learn the dynamics of intra-class and inter-class features within the intra-view.
+To address cross-view discrepancies and inconsistencies, we propose a novel information consistency evolution learning (ICEL) module which combines neighborhood consistency and mutual information optimization to learn consistent and discriminative cross-view representations, bridging the discriminability in intra-view learned by the DHML module to form a synergistic learning system.
+Extensive experiments demonstrate that our method, which employs a lightweight backbone, outperforms existing self-supervised methods on three DVGL benchmarks and even surpasses several state-of-the-art supervised methods.
 The remainder of this paper is organized as follows.
-Section II systematically reviews prior research.
-In Section III, the proposed self-supervised UVGL method is presented in detail.
+Section II systematically reviews previous research.
+In Section III, the proposed self-supervised learning method is presented in detail.
 The experimental results are reported and analyzed in Section IV.
@@
 II Related Works
-This section establishes a systematic and comprehensive theoretical framework to support the development of the proposed self-supervised UVGL method.
-Although this work focuses on UVGL as a sub-task of cross-view geo-localization (CVGL), we first review the overall progress in supervised CVGL.
-We then discuss recent advances in self-supervised CVGL methods.
-Finally, we summarize main methodes in self-supervised ReID.
+This section establishes a systematic and comprehensive theoretical framework to support the development of the proposed self-supervised learning method for DVGL.
+DVGL is typically regarded as a sub-task of cross-view geo-localization (CVGL).
+To provide a more comprehensive understanding of the field, we first review the overall progress in CVGL, including both supervised and self-supervised methods.
+Then, we summarize the main methods in self-supervised ReID that have provided analogous problem-solving paradigms.
 II-A Supervised Cross-View Geo-localization
 CVGL aims to achieve high-precision localization of query images by retrieving GPS-tagged satellite images.
-Early research in CVGL primarily focused on the ground-view CVGL [19, 20, 21], leading to the development of several benchmark datasets, such as CVUSA [22], CVACT [23], and VIGOR [24].
+Early research focused mainly on the CVGL in ground-view [19] with several benchmark datasets, such as CVUSA [20], CVACT [21], and VIGOR [22].
 To address the challenge of ground-view CVGL, researchers have proposed various methods.
-CVM-Net [25] employed NetVLAD [26] to encode images as global descriptors, thereby reducing visual discrepancies caused by viewpoint variations.
+CVM-Net [23] employed NetVLAD [24] to encode images as global descriptors to reduce visual discrepancies caused by viewpoint variations.
 TransGeo [25] utilized attention-based zooming and sharpness-aware optimization to enhance detail learning and improve model generalization.
@@
 Sample4Geo [28] introduced a hard negative mining strategy, focusing on challenging negative samples to further improve discriminative capability.
-With the continuous development and application of drone technology, UVGL has gradually become an important research direction and achieves significant progress [30, 31, 32].
-Several benchmark datasets for UVGL have been proposed, such as University-1652[12], SUES-200, [33] and DenseUAV [34], which have driven further research in this field.
-In this scenario, a variety of methods have been proposed by researchers.
-LPN [35] mined fine-grained features through local pattern partitioning, improving the ability to capture local details.
-IFS [36] utilized a multi-branch strategy to effectively fuse global and local features, enhancing the diversity and robustness of feature representations.
-CAMP [30] improved cross-view matching accuracy by contrastive attribute mining and position-aware partitioning, while DAC [4] enhanced feature consistency and matching accuracy by introducing domain alignment and scene consistency constraints.
-MEAN [2] adopted progressive embedding diversification, global-to-local associations, and cross-domain enhanced alignment to further improve feature representation and alignment.
-Nevertheless, the remarkable performances exhibited by these UVGL methods require extensive pre-processed UAV-satellite image pairs.
-In this work, we pivot our focus to the realm of self-supervised UVGL, where prior pairing relationships are absent, presenting important applications for real-world UVGL deployments.
+With the continuous development and application of drone technology, DVGL has gradually become an important research direction and has achieved significant progress [29, 30, 31].
+Several benchmark datasets for DVGL have been proposed, such as University-1652 [12], SUES-200 [32], and DenseUAV [33], which have driven further research in this field.
+In this scenario, researchers have proposed a variety of methods.
+LPN [34] extracted fine-grained features through local pattern partitioning to improve the ability to capture local details.
+IFS [35] utilized a multi-branch strategy to effectively fuse global and local features.
+CAMP [29] improved the accuracy of cross-view matching by contrastive attribute mining and position-aware partitioning, while DAC [3] enhanced feature consistency and matching accuracy by introducing domain alignment and scene consistency constraints.
+MEAN [2] adopted progressive embedding diversification, global-to-local associations, and cross-domain enhanced alignment to further improve the representation and alignment of features.
+Nevertheless, the remarkable performances exhibited by these DVGL methods require extensive pre-processed drone-satellite image pairs.
+In this work, we concentrate on self-supervised learning approaches for DVGL, particularly addressing scenarios where labeled pairings are unavailable, which is crucial for real-world applications.
 II-B Self-Supervised Cross-View Geo-localization
-Due to the dependence of CVGL on extensive pre-processed image pairs, which significantly increases costs and limits applications for real-world, researchers have begun to explore self-supervised CVGL methods.
-The work in [37] proposed a self-supervised method that leveraged correspondence-free projection to transform ground panorama images into bird’s-eye view images.
+Due to the dependence of CVGL on extensive pre-processed image pairs, which significantly increases costs and limits applications for the real world, researchers have begun to explore self-supervised learning methods for CVGL.
+The work in [36] proposed a self-supervised learning method that used correspondence-free projection to transform ground panorama images into bird’s eye view images.
 It also applied CycleGAN [37] to generate fake satellite images, coupled with an additional SAFA module [38] to reduce discrepancies.
 However, this explicit alignment method has inherent limitations in capturing the deep semantic relationships of cross-view features and may introduce noise.
-Furthermore, the non-end-to-end training framework employed significantly increases training time and computational costs, reducing the efficiency and practical applicability.
-Additionally, existing work [5] utilized frozen foundation models (FMs) [40, 41] to extract features, and a self-supervised adapter was introduced to mitigate visual discrepancies between different point of view.
+Furthermore, the non-end-to-end training method employed significantly increases training time and computational costs with low practical applicability.
+Additionally, existing work [4] utilized frozen foundation models (FMs) [39] to extract features, and a self-supervised adapter was introduced to mitigate visual discrepancies between different points of view.
 However, this method heavily relied on the ability of FMs for feature extraction.
 When FMs fail to effectively handle viewpoint discrepancies, the performance of the adapter is also affected.
-To reduce reliance on the feature extractor, CDIKTNet [42] trained with a small amount of paired UAV-satellite images to optimize the initial feature distribution, and then performed clustering to generate pseudo-labels, thereby improving the quality of pseudo-labels generation.
-However, this method inherently depends on the feature extractor, as it still requires limited paired UAV-satellite images for effective learning.
+To reduce reliance on the feature extractor, CDIKTNet [40] trained with a small amount of paired drone-satellite images to optimize the initial feature distribution and then performed clustering to generate high-quality pseudo-labels.
+However, this method inherently depends on the feature extractor, as it still requires limited paired drone-satellite images for effective learning.
 The above methods have made meaningful exploratory attempts.
@@
 In contrast, our method is fully end-to-end, eliminating the need for additional training processes and completely removing the reliance on feature extractor initialization.
-II-C Self-Supervised Person Re-Identification
-ReID [43, 44] aims at matching the same person image captured by non-overlapping cameras.
+II-C Methodological Parallels from Self-Supervised ReID
+ReID [41] aims to match the same person image captured by non-overlapping cameras.
 Due to its critical role in video surveillance, ReID has attracted extensive research attention and achieved remarkable progress.
+It has provided a fully end-to-end framework to eliminate the need for additional training processes and completely remove the reliance on feature extractor initialization, which offers strong inspiration for our consequent work in the DVGL task.
 However, these achievements are facilitated by extensive human-labeled data.
@@
 Among them, clustering algorithms are widely employed to generate pseudo-labels, which serve as supervision signals for model training.
-In addition, benefiting from the introduction of instance-level contrastive learning [45] with memory banks and momentum update strategies [46], many methods [47, 48] consider each unlabeled sample as an individual class to learn discriminative instance-level feature representations.
-However, these instance-level features was modified largely during training, resulting in unstable representations [49].
+In addition, benefitting from the introduction of instance-level contrastive learning [42] with memory banks and momentum update strategies [43], many methods [44, 45] consider each unlabeled sample as an individual class to learn discriminative instance-level feature representations.
+However, these instance-level features were largely modified during training, which resulted in unstable representations [46].
 To address this issue, recent studies [47, 48] have begun exploring cluster-level relationships to discover more stable and robust associations among pedestrian representations.
-Cluster Contrast [17] addresses the problem of cluster inconsistency by performing clustering and introducing a distinct representation for each cluster.
+Cluster-Contrast [17] contrasts instance features with these cluster-level representations to address the issue of inconsistent instance discrimination.
 Nevertheless, in cross-modal ReID scenarios, where images are captured by heterogeneous sensors, modality discrepancies hinder the effectiveness of these methods in learning modality-invariant features.
 ADCA [18] introduced a dual-path contrastive framework to bridge modality gaps and enhance cross-modal feature alignment.
-Building on this, SDCL [52] employs a dual-path Transformer for shallow-to-deep contrastive learning, and refines pseudo-labels via consistency constraints across feature hierarchies, leading to more stable and discriminative cross-modal representations.
-The above methods focus on alleviating the dependency on human-labeled data in the ReID and provide valuable insights for UVGL.
-However, UVGL suffers from more severe viewpoint and scale variations.
-Furthermore, the spatial continuity of geographic environments inevitably induces feature ambiguity, severely degrading clustering effectiveness.
-Thereby, we incorporate instance-level and cluster-level features into a self-supervised UVGL method to learn more discriminative and robust cross-view representations.
-III METHOD
+Building on this, SDCL [49] employed a dual-path Transformer for shallow-to-deep contrastive learning, and refined pseudo-labels via consistency constraints across feature hierarchies to achieve more stable and discriminative cross-modal representations.
+The above methods focus on alleviating the dependency on human-labeled data in the ReID and provide valuable insights for DVGL.
+However, DVGL suffers from more severe viewpoint and scale variations.
+Furthermore, the spatial continuity of geographic environments inevitably induces feature ambiguity, which would severely degrade clustering effectiveness.
+A cluster-based self-supervised framework, which is developed concurrently with our work, is proposed with a similar ReID framework that employs intermediate state transition and local-to-global matching for cross-view alignment to mitigate viewpoint discrepancies.
+However, this method relies on explicit appearance-level transformations rather than intrinsically modeling cross-view semantic features, and lacks end-to-end optimization and uncertainty-aware pseudo-label refinement, making it sensitive to noise and limiting its adaptability.
+In this work, we integrate instance-level and cluster-level features in a self-supervised method for DVGL by employing reliable neighbor learning to facilitate implicit cross-view alignment and improve the discriminative power and robustness of the learned representations.
+III End-to-End Self-Supervised Learning Method
 As illustrated in Fig.
-3, the proposed framework is composed of two key modules, DHML and ICEL.
+3, the proposed method is composed of two key modules, DHML and ICEL.
 Additionally, a PLE strategy is employed to optimize the self-supervised learning process.
 Specifically, we adopt a weight-shared ConvNeXt-Tiny network as the backbone, equipped with cluster memory and a dual-path contrastive learning strategy to learn intra-view feature representations.
-To capture the dynamic variations of intra-class features caused by changes in viewpoint and scale, the DHML module establishes a collaborative relationship between short-term and long-term memory, thereby enhancing the discriminability of intra-view feature representations.
+To capture the dynamic variations of intra-class features caused by changes in viewpoint and scale, the DHML module establishes a collaborative relationship between short-term and long-term memory.
+DHML module enables efficient discriminability of intra-view feature representations.
 Furthermore, the ICEL module leverages instance memory to construct neighborhood-driven dynamic constraints.
-By integrating neighborhood consistency with mutual information optimization, it guides cross-view features to form unified and consistent representations.
-Problem Formulation: In the UVGL, we are given a set of paired UAV-satellite images, denoted by .
-Here, represents UAV-view images and represents satellite-view images.
-Unlike the supervised setting, our self-supervised method is developed without using paired UAV-satellite images, where the correspondences between UAV-view and satellite-view images are entirely unavailable.
-Thereby, we randomly sample UAV and satellite images independently during training.
-The objective is to discover reliable cross-view correspondences under such conditions and accomplish the UVGL task.
+By integrating neighborhood consistency with mutual information optimization, it guides cross-view features to form consistent and unified representations.
+Problem Formulation: In DVGL, a set of paired drone-satellite images is denoted as , where represents the drone-view image and represents the satellite-view image.
+Unlike the supervised setting, our self-supervised method is developed without using paired drone-satellite images, where the correspondences between drone-view and satellite-view images are entirely unavailable.
+The objective is to discover reliable cross-view correspondences under such conditions and to accomplish the DVGL task.
 III-A Dual-Path Contrastive Learning Baseline
-Cluster-based contrastive learning methods have achieved remarkable success in various fields.
-However, in UVGL, due to the unique challenges of cross-view scenarios, single-path cluster-based contrastive learning struggles to effectively capture view-specific features.
-To address this issue, we design the baseline network with a dual-path architecture, where UAV-view and satellite-view features are extracted using a shared backbone and clustered to assign pseudo-labels.
+Cluster-based contrastive learning methods have achieved remarkable success in various fields [17, 50].
+However, in DVGL, the unique challenges of cross-view scenarios, as discussed in Section I, make it difficult for single-path cluster-based contrastive learning to effectively capture view-specific features.
+Therefore, we design the baseline network with a dual-path architecture, where drone-view and satellite-view features are extracted using a shared backbone and clustered to assign pseudo-labels.
 Subsequently, a dual-path contrastive learning strategy [51] is employed to enhance intra-view representation learning.
-We first introduce the notations used in this paper for better clarity.
+We first introduce the notation for better clarity.
 Let denote a set of drone images with instances, and denote a set of satellite images with instances.
-and represent the feature representations extracted by the shared feature extractor for the drone and satellite views, respectively.
+and represent the feature representations extracted by the shared feature extractor for drone-view and satellite-view, respectively.
 and denote the drone query instance features and satellite query instance features extracted by the feature extractor .
-The processing procedure of the backbone can be described as
-Cross-View Feature Memory Initialization.
-At the beginning of each training epoch, the feature memories for both views are initialized.
-Specifically, the cluster representations of the drone view and the satellite view are stored in the drone memory dictionary and satellite memory dictionary , respectively, for view-specific memory initialization.
+Arbitrary Sampling and Backbone-based Feature Extraction.
+To simulate real-world scenarios where drone and satellite images are not paired, we independently and arbitrarily sample drone and satellite images at each training epoch to construct unpaired batches.
+Each image is subsequently processed by a shared backbone network to extract features.
+The backbone processing procedure can be described as
+Cross-View Cluster Memory Initialization.
+At the beginning of each training epoch, the feature memories for both views are initialized into clusters and clusters, respectively.
+Specifically, the cluster centroids of the drone-view and the satellite-view are stored in the drone memory dictionary and the satellite memory dictionary , respectively, for view-specific memory initialization.
 This process can be formalized as
-where , , , , denotes the -th cluster set in drone or satellite view, indicates the number of instances per cluster.
-Cross-View Feature Memory Updating.
-During the training phase, we randomly sample distinct localization points from the training set, with each point comprising a fixed number of instances.
+where denotes the -th cluster set in drone-view, , and denotes the -th cluster set in satellite-view, . indicates the number of instances per cluster.
+Cross-View Cluster Memory Updating.
+During the training phase, we randomly sample different localization points from the training set, with each point comprising a fixed number of instances.
 This sampling strategy results in a minibatch containing a total of query images.
 Subsequently, the query features extracted from the minibatch are integrated into the cluster representations via a momentum-based update mechanism.
-Through this method, we iteratively update the features for both perspectives.
-This process can be formalized as
-where is the momentum updating factor. is the iteration number.
-Cross-View Contrastive Loss. Given drone and satellite query and , we compute the contrastive loss for drone view and satellite view by the following equations:
+The cluster centroids and are updated iteratively as
+where is the momentum updating factor.
+Cross-View Cluster Memory Contrastive Loss. Given drone and satellite query and , we compute the contrastive loss for drone-view and satellite-view as
 where denotes the temperature hyper-parameter.
-The cluster centroids are utilized as feature vectors at the cluster level to compute the distances between the query instance and all clusters.
-Here, corresponds to the positive cluster feature associated with .
-The final optimization for baseline is denoted by the following combination:
+Cluster centroids and are utilized as feature vectors at the cluster level to compute the distances between the query instance and all clusters.
+Here, and are positive cluster features associated with and .
+The final optimization for baseline is denoted as
 III-B Dynamic Hierarchical Memory Learning
-Geographic environments inherently exhibit spatial continuity, where neighboring regions usually present similar visual patterns, while distant regions display significant differences.
+Geographic environments inherently exhibit spatial continuity, where neighboring regions usually present similar visual patterns, while distant regions show significant differences.
 However, such spatial characteristics often introduce ambiguity in intra-view feature representations.
-To address this, we propose a DHML module that jointly exploits long-term and short-term memory to capture spatial dynamics and enhance the discriminability of intra-view features.
-Consistent with the baseline network, we construct the dynamic hierarchical memory learning in a dual-path manner, using drone feature extraction as an illustrative example.
-To build the long-term memory, we first construct two novel clustering memory dictionaries according to Eq.2, which are used to store the short-term and long-term memories, respectively.
-where, represent the short-term memory embeddings that encode the cluster representations of the drone view, and represent the long-term memory embeddings that store historically accumulated stable representations of the drone view.
-Unlike the weight updating scheme in Eq.4 and Eq.5, we assign different weights to the input features and the historical memory during the long-term memory update.
+To address this issue, we propose a DHML module that jointly exploits long-term and short-term memory to capture spatial dynamics and enhance the discriminability of intra-view features.
+Consistent with the baseline network, we construct dynamic hierarchical memory learning in a dual-path manner, using drone feature extraction as an illustrative example.
+To build long-term memory, we first construct two novel clustering memory dictionaries according to Eq.(2), which are used to store short-term and long-term memories, respectively.
+They are expressed as
+where represent the short-term memory embeddings that encode the cluster representations of the drone-view, and represent the long-term memory embeddings that store historically accumulated stable representations of the drone-view.
+Unlike the weight updating scheme in Eq.(4), we assign different weights to the input features and historical memory during the long-term memory update.
 Specifically, since the momentum parameter is set to 0.2 throughout this study, we enhance the influence of historical memory during the initialization of long-term memory updates.
 The update procedure can defined as
-where, represents the drone query instance features. This obtain a new dictionary of long-term memory .
-Additional, we further introduce an adaptive weight updating mechanism to update the long-term memory.
+where represents the drone query instance features. This obtains a new dictionary of long-term memory.
+In addition, we further introduce an adaptive weight update mechanism to update the long-term memory.
 In detail, we compute the euclidean distance between the current input and its corresponding long-term memory, then take the average over all samples in the batch.
 This average distance is then mapped to the range via the sigmoid function, yielding an adaptive update coefficient:
-where, denotes the batch size and represents the sigmoid function. The short-term memory is subsequently updated as follows:
+where denotes the batch size and represents the sigmoid function. The short-term memory is subsequently updated as
 Then, the long-term and short-term memories are fused in a weighted manner to yield the combined feature representation:
@@
 where denotes the temperature hyper-parameter.
-The combined cluster centroids are utilized as feature vectors at the cluster level to compute the distances between the query instance and all clusters.
+The combined cluster centroids are used as feature vectors at the cluster level to calculate the distances between the query instance and all clusters.
 Here, corresponds to the positive cluster feature associated with .
-Finally, we introduce a fusion factor to perform weighted fusion of and , aiming to enhance the discriminability of intra-view feature representations and obtain the optimized loss .
+Finally, we introduce a fusion factor to perform a weighted fusion of and , aiming to enhance the discriminability of intra-view feature representations and obtain the optimized loss .
 Similarly, can be obtained in the same way.
@@
 Therefore, the proposed ICEL module employs a threshold-based and ranking-based neighborhood selection strategy to achieve effective cross-view feature alignment.
-By promoting the consistency evolution of cross-view feature representations, ICEL enhances both the discriminability and representation capability of cross-view features.
+By promoting the consistency evolution of cross-view feature representations, ICEL enhances both the discriminability and the representation capability of cross-view features.
 In the ICEL module, unlike the baseline and DHML module, we employ instance-level features instead of cluster-level representations.
@@
 This process can be described as
-where and denote the drone and satellite instance features stored in memory dictionaries and , respectively.
-In addition, we define the similarity between a given drone (or satellite) query and each drone (or satellite) instance in the training set as
-where and are the indices of queries and instances. Thus, we obtain intra-view similarity , and cross-view similarity , .
+where , , and , , denote the drone and satellite instance features stored in memory dictionaries and , respectively.
+To facilitate the following description, we take the scenario where the drone-view serves as the query as an illustrative example.
+The similarity between a given drone query that denotes the -th query sample in the drone-view, .
+denote the batch size of the query .
+Each drone instance and each satellite instance in the training set are compared to drone query by computing a similarity score, respectively.
+This is defined as
 Then, we adopt a threshold-based neighborhood selection strategy to select reliable neighbors, which can be defined as
-where represents the set of drone feature instances selected as valid intra-view neighbors.
-denotes the set of drone feature instances that exhibit strong cross-view associations with satellite features.
+where represents the set of drone feature instances selected as valid intra-view neighbor, and denotes the set of drone feature instances that exhibit strong cross-view associations with satellite features.
 is a predefined threshold.
-The and can be obtained by the same way.
 Additionally, we introduce a ranking-based neighborhood selection strategy.
@@
 The process can be defined as
-where , and , denote the selected intra-view and cross-view neighborhood sets, respectively, with and containing exactly elements, and and containing exactly elements.
-Similarly, the neighborhood sets , , , and can be obtained following the same selection strategy.
+where and denote the selected intra-view and cross-view neighborhood sets, respectively. and can be obtained following the same selection strategy.
 Based on the constructed neighborhoods, the normalized similarity distribution is computed within the larger neighborhood to promote the aggregation of similar samples in both intra-view and cross-view scenarios.
@@
 The intra-view and cross-view losses can be formulated as
-where represents the batch size of the query , and denote the total number of neighborhood samples within the set and .
+where and denote the total number of neighborhood samples within the set and .
 Incorporating neighborhoods , and , , we introduce consistency loss to enhance alignment of relative similarity distributions within the large-scale neighborhood across both intra-view and cross-view representations.
 The intra-view loss and cross-view loss under this consistency loss can be formulated as
-where is the softmax-normalized similarity distribution, is used for summation to normalize the similarity scores over all neighborhood samples. and can be obtained in the same way.
+where and are the softmax-normalized similarity distribution.
 Within a more restrictive and smaller set of neighbors, a mutual information constraint is proposed to learn intra-view and cross-view features representations.
-This complements the consistency constraint described above, achieving a balance between global structural alignment and local discriminative information.
-Within the selected neighborhood, we first compute, for each neighbor sample (where ), the similarity between the deep query feature and the neighbor feature using the similarity function .
-Based on these similarity scores, the posterior probability for each neighbor is defined via a softmax normalization as:
-where denote the total number of neighborhood samples within the set and .
-Therefore, the intra-view and cross-view mutual information losses and can be derived and are defined as follows:
-where and represent the normalized similarity distributions over the selected intra-view and cross-view neighbors, respectively. and can be obtain in the same way.
-For each view combination, the loss is computed as the weighted sum of three components: the original neighborhood loss , the strict neighborhood loss , and the extended neighborhood loss :
-where and denote the drone or satellite view.
-The total neighbor learning loss is then formulated as
+This complements the consistency constraint described above and achieves a balance between global structural alignment and local discriminative information.
+Within the selected neighbors, we compute the similarity between the query and each neighbor and using the function , and apply softmax to obtain the posterior probability for each neighbor.
+It can be described as
+Therefore, the intra-view and cross-view mutual information losses and can be derived and are defined as
+For each view, the loss is formulated as a weighted sum of the original neighborhood loss , the strict neighborhood loss , and the extended neighborhood loss .
+Specifically, the intra-view loss and the cross-view loss are respectively defined as
+where and are hyperparameters controlling the contribution of strict and extended neighborhood components.
+When the satellate view as the query, the and can be obtained in the same way.
+The total neighbor learning loss is computed as .
 Finally, the overall loss of our method is defined as
@@
 In self-supervised learning, the quality of pseudo-labels exerts a decisive influence on model performance.
-Inspired by [52], our proposes a pseudo-label optimization strategy based on dynamic feature robustness.
+Inspired by [49], we propose a pseudo-label optimization strategy based on dynamic feature robustness.
 Specifically, we introduce small perturbations to the original feature representations and exploit the top-K neighbor consistency between the original and perturbed features to identify high-confidence samples while filtering out noisy instances.
-Then, a spatial smoothing operation based on the similarity matrix is applied to refine the pseudo-labels, enforcing local label consistency.
-We first employe small Gaussian noise into the original satellite instances feature and the original drone instances features to obtain perturbed features, which are used for more robust similarity estimation and subsequent cluster consistency evaluation.
+Then, a spatial smoothing operation based on the similarity matrix is applied to refine the pseudo-labels to enhance local label consistency.
+Small Gaussian noise is used in the initial satellite instances feature and the initial drone instances features to obtain perturbed features, where the denotes the standard deviation.
 It can be formulated as
-where, with the similarities and , we can get two cross-view ranking lists and for and , which are denoted as and .
-The label of the -th similar infrared instance in two ranking lists can be represented as
-where and are refined -th cross-view labels of and through the ranking on cross-view initial features and perturbed features similarities.
-So, we associate the cross-modality labels with the intersection of two label sets to investigate collaborative ranking consistency. This process can be described as
-where records the samples with the same identity of top- identity in instances from top-1 to top-.
-The reliable refined cross-view label of can be expressed by the label of maximum count in .
-where denotes the counting function. The operation traverses and finds the labels with maximum number as cross-view refined labels.
-Then, we convert the label list to the form of one-hot code matrix by setting the column according to the refined labels to 1 and the rest to 0.Then,two initial and perturbed homogeneous initial and deep similarity matrices are constructed to investigate intra-view ranking consistency, enhancing the precision of refined cross-view pseudo labels by:
-where represents the intra-view similarity structure.
-In order to explore the intra-view initial and perturbed collaboration, we calculate the sum of and by:
-where indicates the consistency of initial and perturbed similarity matrix.
-We keep the 5-max values of in each row to 1 and set the rest to 0, acquiring the ranking relations. The process of intra-modality ranking smoothing is formulated as follows:
-where is the final refined cross-view label matrix of satellite instance. In , the column number of the maximum value in each row is the refined label of samples.
+Using the definition of the similarity score shown in Eq.(16), two cross-view ranking lists can be obtained, which are denoted as and , .
+The label of the -th similar drone instance in two ranking lists can be denoted as and .
+Next, the cross-view labels are associated with the intersection of two label sets to investigate collaborative ranking consistency.
+This process can be described as
+where records the samples with the same identity in instances.
+Accordingly, we update the cross-view pseudo-label for using the label with the highest frequency in , which reflects the most stable identity under both original and perturbed feature rankings.
+This final refined label is denoted as .
+Then, we convert the label list into the form of one-hot code matrix by setting the column according to the refined labels to 1 and the rest to 0.
+In order to explore the intra-view initial and perturbed collaboration, the initial and perturbed similarity matrices and are calculated, and their sum is used to investigate intra-view ranking consistency.
+Then we keep the 5-max values of in each row at 1 and set the rest to 0, acquiring the ranking relations.
+The process of intra-view ranking smoothing is formulated as
+where is the final refined cross-view label matrix of satellite instance. In , the column number of the maximum value in each row is the refined label of the samples.
 University-1652 | |||||||
 DroneSatellite | SatelliteDrone | ||||||
-| Model | Venue | Supervision Ratio | Paramars(M) | R@1 | AP | R@1 | AP |
+Model | Venue | Supervision Ratio | Parameters(M) | R@1 | AP | R@1 | AP |
 MuSe-Net[52] | PR’2024 | 100(%) | 50.47 | 74.48 | 77.83 | 88.02 | 75.10 |
@@
 CDIKTNet[40] | arXiv’2025 | 0(%) | 91.52 | 83.30 | 85.73 | 87.73 | 76.53 |
+Wang[9] | AAAI’2025 | 0(%) | - | 85.95 | 90.33 | 94.01 | 82.66 |
 Ours | - | 0(%) | 28.59 | 90.17 | 91.67 | 95.01 | 88.95 |
@@
 IV-A Experimental Datasets and Evaluation Metrics
-To evaluate our self-supervised method, we conduct experiments on three benchmark datasets including University-1652 [12], SUES-200 [33] and DenseUAV [54].
-University-1652 contains images from 1,652 buildings across 72 universities worldwide.
-It includes drone, satellite, and ground-level views.
-The training set consists of 701 buildings from 33 universities, and the test set covers 951 buildings from 39 geographically disjoint universities.
-SUES-200 focuses on altitude diversity in aerial imagery, containing 200 locations with 120 for training and 80 for testing.
-Each location provides one satellite image and aerial images captured at four altitudes (150m, 200m, 250m, and 300m), covering urban areas, natural landscapes, and water bodies.
-SUES-200 evaluates cross-view retrieval under varying flight heights, reflecting real-world UAV scenarios.
-DenseUAV focuses on UAV self-positioning in low-altitude urban environments, containing over 27,000 UAV and satellite images collected from 14 university campuses.
-It features dense sampling, multi-scale satellite images, and varying altitudes (80m, 90m, and 100m) under multiple temporal conditions.
-The training set includes 6,768 UAV images and 13,536 satellite images, while the test set comprises 2,331 UAV images and 4,662 satellite images from non-overlapping areas.
+To evaluate the proposed method, we conduct experiments on three benchmark datasets, including University-1652 [12], SUES-200 [32] and DenseUAV [52].
+University-1652 contains abundant images of 1,652 buildings in 72 universities around the world.
+It involves drone, satellite, and ground-level views.
+The training set consists of 701 buildings from 33 universities and the test set covers 951 buildings from 39 geographically nonaffiliated universities.
+SUES-200 focuses on altitude diversity in aerial imagery, containing 200 locations, with 120 designated for training and 80 for testing.
+Each location in the dataset contributes one satellite image along with aerial images captured at four distinct altitudes, which involve urban areas, natural landscapes, and water bodies.
+SUES-200 evaluates cross-view retrieval under varying flight heights, reflecting real-world drone scenarios.
+DenseUAV focuses on drone self-positioning in low-altitude urban environments with more than 27,000 drone and satellite images collected from 14 university campuses.
+It is characterized by dense sampling, multi-scale satellite imagery, and diverse altitudes captured across multiple temporal conditions.
 DenseUAV poses challenges such as cross-view matching, neighbor confusion, and spatial-temporal variation.
-We evaluate the models using Recall@K (R@K) and Average Precision (AP), where R@K measures the proportion of correct matches among the Top-results and AP balances precision and recall.
-Model efficiency is assessed by the number of parameters, reflecting portability under resource constraints.
+We evaluate the models using Recall@K (R@K) and Average Precision (AP), where R@K measures the proportion of correct matches among the Top-results and AP evaluates the overall precision-recall performance.
+Model efficiency is assessed by the number of parameters to reflect portability under resource constraints.
 All comparisons are conducted using the best configurations for each method.
@@
 During both training and testing, all input images are resized to .
-We employ several data augmentation techniques, including random cropping, random horizontal flipping, and random rotation.
-The batch size is set to 64, and DBSCAN [8] is used to generate pseudo labels.
+We employ several data-augmentation techniques, including random cropping, random horizontal flipping, and random rotation.
+The batch size is set to 64, and DBSCAN [7] is used to generate pseudo-labels.
 For optimization, we use the SGD optimizer with an initial learning rate of 0.001, training the model for 30 epochs in total.
 All experiments are conducted under the PyTorch framework on an Ubuntu 22.04 system equipped with four NVIDIA RTX 4090 GPUs.
-Additionally, in the University-1652 and SUES-200 datasets,we replicate each satellite feature 50 times for clustering purposes, leaving UAV features unchanged.
-In DenseUAV, we replicate UAV features four times at each altitude, and replicate the satellite features three times for each time point and altitude.
+Additionally, for the University-1652 and SUES-200 datasets, we apply feature replication only on the training sets by duplicating each satellite feature 50 times for clustering purposes, while drone features remain unchanged.
+The test sets are kept intact without any modification.
+In the DenseUAV dataset, drone features are replicated four times at each altitude, and satellite features are replicated three times for each time point and altitude exclusively in the training set; the test set remains unchanged.
 IV-C Comparison with State-of-the-art Methods
-To evaluate the effectiveness of our method, we compare it with the state-of-the-art supervised methods MuSe-Net [54], MCCG [1], SDPL [3], CCR [32], Sample4Geo [29], SRLN[31], MEAN [2] and DAC [4], and the recent self-supervised methods EM-CVGL [5], CDIKTNet [42].
+To evaluate the effectiveness of our method, we compare it with the state-of-the-art supervised methods MuSe-Net [52], MCCG [1], SDPL [53], CCR [31], Sample4Geo [28], SRLN [30], MEAN [2] and DAC [3], and the self-supervised methods EM-CVGL [4], CDIKTNet [40] and Wang [9].
 Results on University-1652.
-As shown in Table I, our method exhibits superior performance compared with the current advanced self-supervised methods.
+As shown in Table I, our method exhibits a superior performance compared to current advanced self-supervised methods.
 Specifically, our method achieves R@1/AP scores of 90.17%/91.67% and 95.01%/88.95% under the DroneSatellite and SatelliteDrone settings, respectively.
-Compared with EM-CVGL, our method achieves significant improvements, outperforming it by approximately 20% in R@1 and 17% in AP on the DroneSatellite setting, and by 16% in R@1 and 28% in AP on the SatelliteDrone setting.
-Furthermore, even compared with CDIKTNet, which utilizes paired data from other domains for self-supervised fine-tuning, our method still consistently achieves superior performance without relying on any cross-domain paired data.
-We further benchmark against eight representative supervised methods, and the results show that our method clearly surpasses some of them, including MuSe-Net, MCCG, and SDPL.
+Compared to EM-CVGL, our method achieves significant improvements, outperforming it by approximately 20% in R@1 and 17% in AP on the DroneSatellite setting, and by 16% in R@1 and 28% in AP on the SatelliteDrone setting.
+Furthermore, even compared to CDIKTNet, which utilizes paired data from other domains for self-supervised fine-tuning, our method still consistently achieves superior performance without relying on any cross-domain paired data.
+Furthermore, compared with the latest clustering-based self-supervised methods, our approach still achieves significant improvements.
+We further benchmark with eight representative supervised methods, and the results show that our method clearly surpasses some of them, including MuSe-Net, MCCG, and SDPL.
 Furthermore, with self-supervised pre-training and subsequent supervised fine-tuning, our method achieves competitive performance under the supervised setting, comparable to CCR, Sample4Geo, and SRLN.
@@
 150m | 200m | 250m | 300m | ||||||||
-| Model | Venue | Supervision Ratio | Paramars(M) | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
+Model | Venue | Supervision Ratio | Parameters(M) | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
 MCCG[1] | TCSVT’2023 | 100(%) | 56.65 | 82.22 | 85.47 | 89.38 | 91.41 | 93.82 | 95.04 | 95.07 | 96.20 |
@@
 CDIKTNet[40] | arXiv’2025 | 0(%) | 91.52 | 82.75 | 85.25 | 89.35 | 91.13 | 93.15 | 94.39 | 95.18 | 96.12 |
+Wang[9] | AAAI’2025 | 0(%) | - | 76.90 | 84.95 | 87.88 | 92.60 | 92.98 | 95.66 | 95.10 | 96.92 |
 Ours† | - | 0(%) | 28.59 | 82.60 | 85.60 | 90.48 | 92.28 | 94.33 | 95.44 | 96.78 | 97.42 |
@@
 150m | 200m | 250m | 300m | ||||||||
-| Model | Venue | Supervision Ratio | Paramars(M) | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
+Model | Venue | Supervision Ratio | Parameters(M) | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
 CCR[31] | TCSVT’2024 | 100(%) | 156.57 | 92.50 | 88.54 | 97.50 | 95.22 | 97.50 | 97.10 | 97.50 | 97.49 |
@@
 CDIKTNet[40] | arXiv’2025 | 0(%) | 91.52 | 88.75 | 80.33 | 93.75 | 88.17 | 95.00 | 92.15 | 98.75 | 94.37 |
+Wang[9] | AAAI’2025 | 0(%) | - | 87.50 | 74.81 | 92.50 | 87.15 | 96.25 | 91.20 | 98.75 | 94.52 |
 Ours† | - | 0(%) | 28.59 | 91.25 | 81.28 | 97.50 | 90.62 | 98.75 | 94.74 | 97.50 | 96.17 |
@@
 Results on SUES-200.
-As shown in Table II, As shown in Table II, our method, when directly transferring the pre-trained model from the University-1652 dataset to the SUES-200 dataset without any re-training, consistently outperforms all existing self-supervised methods and achieves competitive performance compared to supervised methods such as MCCG and SDPL.
+As shown in Table II, our method, when directly transferring the pre-trained model from the University-1652 dataset to the SUES-200 dataset without any re-training, consistently outperforms all existing self-supervised methods and achieves competitive performance compared to supervised methods such as MCCG and SDPL.
 Moreover, by performing self-supervised re-training on the SUES-200 dataset, our method brings a significant performance boost, achieving results comparable to the supervised CCR.
@@
 These results clearly demonstrate the effectiveness of our method in self-supervised learning, transfer learning, and efficient fine-tuning.
-Results on Cross-Domain Generalization Performance: Our method consistently outperforms existing self-supervised method under cross-domain settings, showing both significant improvements and stable generalization.
-In the Drone to Satellite setting, our model achieves better performance than EM-CVGL at all altitudes.
-At 150 meters, R@1 and AP increase by approximately 22% and 20%; at 200 meters, the gains are around 19% and 16%; at 250 meters, nearly 14% and 12%; and even at 300 meters, our method still outperforms by about 11% and 9%.
+Results on Cross-Domain Generalization Performance.
+As shown in Table III, our method consistently outperforms existing self-supervised methods in cross-domain settings, showing both significant improvements and stable generalization.
+Specifically, in the Drone to Satellite setting, our model achieves better performance than EM-CVGL at all altitudes.
+At 150 meters, R@1 and AP increase by approximately 22% and 20%.
+At 200 meters, the gains are around 19% and 16%; at 250 meters, nearly 14% and 12%.
+Even at 300 meters, our method still outperforms by about 11% and 9%.
 In the Satellite to Drone scenario, similar improvements are observed.
 At 150 meters, R@1 and AP improve by approximately 18% and 25%; at 200 meters, by 10% and 20%; at 250 meters, by 6% and 13%; and at 300 meters, by 3% and 10%.
-These results confirm the robustness and strong cross-domain generalization capability of our method across varying altitudes.Remarkably, our method even surpasses all existing state-of-the-art supervised methods in terms of cross-domain generalization.
+These results confirm the robustness and strong cross-domain generalization capability of our method across varying altitudes.
+Remarkably, our method even surpasses all existing state-of-the-art supervised methods in terms of cross-domain generalization.
 These results confirm the robustness and strong cross-domain generalization capability of our method across varying altitudes.
@@
 150m | 200m | 250m | 300m | ||||||||
-| Model | Venue | Supervision Ratio | Paramars(M) | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
+Model | Venue | Supervision Ratio | Parameters(M) | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
 MCCG[1] | TCSVT’2023 | 100(%) | 56.65 | 57.62 | 62.80 | 66.83 | 71.60 | 74.25 | 78.35 | 82.55 | 85.27 |
@@
 150m | 200m | 250m | 300m | ||||||||
-| Model | Venue | Supervision Ratio | Paramars(M) | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
+Model | Venue | Supervision Ratio | Parameters(M) | R@1 | AP | R@1 | AP | R@1 | AP | R@1 | AP |
 MCCG[1] | TCSVT’2023 | 100(%) | 56.65 | 61.25 | 53.51 | 82.50 | 67.06 | 81.25 | 74.99 | 87.50 | 80.20 |
@@
 All height | 80m | 90m | 100m | ||||||||
-| Model | Venue | Supervision Ratio | Paramars(M) | R@1 | R@5 | R@1 | R@5 | R@1 | R@5 | R@1 | R@5 |
+Model | Venue | Supervision Ratio | Parameters(M) | R@1 | R@5 | R@1 | R@5 | R@1 | R@5 | R@1 | R@5 |
 ConvNext-Tiny[33] | TIP’2024 | 100(%) | 30.10 | 60.23 | 81.94 | - | - | - | - | - | - |
@@
 Ours | - | 0(%) | 28.59 | 73.79 | 90.69 | 73.10 | 92.92 | 74.65 | 95.37 | 75.03 | 95.50 |
-Results on DenseUAV: The DenseUAV dataset presents a highly challenging scenario due to its densely sampled nature, where neighboring images exhibit high visual similarity, often leading to severe confusion in UVGL.
-In addition, the dataset includes satellite images captured at different times and UAV images from multiple altitudes, which further increases intra-domain variability and task complexity.
-As shown in Table IV, under the fully unsupervised setting, our method achieves 73.79% R@1 and 90.69% R@5 across all altitudes, significantly outperforming existing self-supervised methods.
-It is noteworthy that the two representative self-supervised baselines consistently yield poor performance on DenseUAV, highlighting their limitations in handling densely sampled and spatiotemporal diverse scenarios.
+Results on DenseUAV.
+The DenseUAV dataset presents a highly challenging scenario due to its densely sampled nature, where neighboring images exhibit high visual similarity, often leading to severe confusion in DVGL.
+In addition, the dataset includes satellite images captured at different times and drone images from multiple altitudes, which further increases intra-domain variability and task complexity.
+As shown in Table IV, in the fully unsupervised setting, our method achieves 73.79% R@1 and 90.69% R@5 across all altitudes, significantly outperforming existing self-supervised methods.
+It is noteworthy that the two representative self-supervised baselines consistently yield poor performance on DenseUAV, due to their limitations in handling densely sampled and spatio-temporal diverse scenarios.
 Furthermore, after supervised fine-tuning, our method surpasses several supervised baselines, including ConvNeXt-Tiny, Sample4Geo, CAMP, and DAC.
@@
 IV-D Ablation Studies
-To thoroughly evaluate the effectiveness of our proposed method, we conduct an ablation study on the University-1652 dataset.
+To thoroughly evaluate the effectiveness of our proposed method, we perform an ablation study on the University-1652 dataset.
 We systematically assess the impact of each individual module and investigate their contributions under different configurations.
@@
 The backbone refers to the untrained ConvNeXt-Tiny network used as a direct feature extractor.
-As shown in Table LABEL:Ablation_Studies, its performance is significantly limited, achieving an R@1 of only 9.55% and 35.24%, and an AP of 11.88% and 13.20% for the Drone Satellite and Satellite Drone, respectively.
+As shown in Table LABEL:Ablation_Studies, its performance is significantly limited, and obtains an R@1 of only 9.55% and 35.24%, and an AP of 11.88% and 13.20% for the Drone Satellite and Satellite Drone, respectively.
 These results indicate that the backbone alone lacks the ability to extract discriminative cross-view features.
 Effectiveness of Dual-Path Contrastive Learning Baseline.
-After incorporating the dual-path contrastive learning baseline, the R@1 improves to 71.06% and 81.31%, while AP reaches 74.91% and 65.26%, respectively.
-Although the baseline primarily learns intra-view feature discrimination, the training process also facilitates a certain degree of alignment between UAV and satellite representations.
+After incorporating the baseline, the R@1 improves to 71.06% and 81.31%, while AP reaches 74.91% and 65.26%, respectively.
+Although the baseline primarily learns intra-view feature discrimination, the training process also facilitates a certain degree of alignment between drone and satellite representations.
 This leads to the formation of a more consistent and unified feature representation space, which serves as a solid foundation for subsequent cross-view learning.
 Effectiveness of Dynamic Hierarchical Memory Learning.
-Compared with the contrastive learning baseline, incorporating DHML leads to significant performance improvements under both settings.
+Compared with the baseline, the incorporation of DHML leads to significant performance improvements in both settings.
 Specifically, R@1 increases by 12.97% and 9.13%, and AP improves by 11.5% and 17.48% , reaching 84.03% and 90.44% in R@1, and 86.41% and 82.74% in AP, respectively.
-These gains can be attributed to the joint modeling of long-term and short-term feature dependencies introduced by DHML, which enhances the capacity to capture spatial structures in UVGL scenarios and enhances intra-view feature representation and discrimination.
+These gains can be attributed to the joint modeling of long-term and short-term feature dependencies introduced by DHML, which enhances the capacity to capture spatial structures in DVGL scenarios and enhances intra-view feature representation and discrimination.
 Setting | University-1652 | |||||||
@@
 Effectiveness of Information Consistency Evolution Learning.
-Although DHML enhances intra-view feature representation and discrimination, its optimization remains limited to within-view information.
+Although DHML enhances intra-view feature representation and discrimination, its optimization remains limited to intra-view information.
 To address this limitation, ICEL is proposed to learn cross-view dependencies.
@@
 As a result, R@1 improves to 86.13% and 51.04%, while AP reaches 88.31% and 82.97%.
-These results further validate the effectiveness of ICEL in enhancing cross-view consistency for UVGL, thereby contributing to overall performance improvement.
-Pseudo-Label Enhancement (PLE): Refining Feature Learning with Robust Labels.
+These results validate the effectiveness of ICEL in enhancing cross-view consistency for DVGL.
+Pseudo-Label Enhancement Refining Feature Learning with Robust Labels.
 To mitigate pseudo-label noise caused by clustering, we apply PLE to improve label reliability.
@@
 This confirms that refining pseudo-label quality significantly enhances the final model’s effectiveness in cross-geographic view matching.
-The ablation study systematically evaluates the contribution of each component in the proposed self-supervised framework.
+The ablation study systematically evaluates the contribution of each component in our method.
 Starting with a backbone that lacks sufficient discriminative power, we first establish a dual-branch contrastive learning baseline.
 On top of this, DHML is introduced to strengthen intra-view feature representation and discriminability, ICEL is employed to facilitate cross-view feature association, and PLE is leveraged for pseudo-label refinement.
-Experimental results show that our method significantly enhances the learning of discriminative and view-consistent representations under self-supervised learning.
+Experimental results show that our method enhances the learning of discriminative and view-consistent representations under self-supervised learning.
 IV-E Visualization
 Feature space and similarity distribution.
-To further illustrate the effectiveness of the our method, We perform feature space (t-SNE[55] map) and similarity distribution visualization for our method, as presented in Fig 4.Comparing (a) Baseline and (b) Ours, it is evident that our method enhances the clustering of cross-view positive samples, effectively bringing the drone and satellite representations of the same location closer together.
-Additionally, the separation between different locations is more distinct, indicating better feature discriminability.
+To further illustrate the effectiveness of the method, we perform a feature space and similarity distribution visualization for our method, as presented in Fig.
+4.
+Comparing (a) Baseline and (b) Ours, it is evident that our method enhances the clustering of cross-view positive samples, effectively bringing the drone and satellite representations of the same location closer together.
+Additionally, the separation between different locations is more distinct that indicates better feature discriminability.
 The second row shows the similarity distribution between drone-satellite positive and negative pairs.
 Our method demonstrates a clearer distinction between positive and negative pairs, with a reduced overlap between distributions, which suggests improved cross-view matching and reduced feature ambiguity.
-This highlights the effectiveness of our method in UVGL.
+This highlights the effectiveness of our method.
 Cluster analysis.
 To validate the efficacy of our method, as shown in Fig.
-5, we illustrate the evolution of cluster counts across epochs for three cross-view datasets.
+5, we illustrate the evolution of cluster counts in each epoch for three datasets.
 In the DenseUAV dataset (2,255 locations), satellite images exhibit significant clustering fluctuations during the early training stages, due to low intra-class discriminability and weak cross-view correlations.
-In contrast, the University-1652 dataset (701 locations) demonstrates greater UAV image clustering instability, while satellite image clusters stabilize more rapidly in the initial epochs.
-The SUES-200 dataset (1,200 locations), with its smaller data size and simpler distribution, exhibits minimal early-stage variation and stabilizes faster than larger datasets.
+In contrast, the University-1652 dataset (701 locations) demonstrates greater drone image clustering instability, while satellite image clusters stabilize more rapidly in the initial epochs.
+The SUES-200 dataset (120 locations), with its smaller data size and simpler distribution, exhibits minimal early-stage variation and stabilizes faster than larger datasets.
 As training progresses, the cluster counts across all datasets gradually converge toward the ground-truth location counts, despite initial inter-class confusion.
-This process validates the ability of our method to progressively learn discriminative representations.
+This process validates the ability of our method to learn discriminative representations progressively.
 Retrieval results.
-To demonstrate the effectiveness of our method, we present retrieval results on the University-1652 and DenseUAV dataset in Fig.
+To demonstrate the effectiveness of our method, we present retrieval results on the University-1652 and DenseUAV datasets in Fig.
 6 and Fig.
@@
 Green borders indicate correct matches, while red borders denote errors, showing that our method significantly improves retrieval accuracy.
-In the University-1652 dataset, the dronesatellite retrieval task presents drone-view images as queries and retrieves corresponding satellite-view images.
-Since each location is associated with a single satellite image, our method consistently achieves Top-1 accuracy, demonstrating its high effectiveness in cross-view matching.
-In the satellitedrone setting, our method accurately retrieves the corresponding drone-view images given satellite queries, further showcasing its strong cross-view localization capability and reliable feature alignment across perspectives.
-In the DenseUAV dataset, the retrieval task becomes significantly more challenging due to dense spatial sampling, multi-scale satellite imagery, and temporal dynamics, all of which exacerbate the difficulty of cross-view representation learning.
-In the Drone Satellite retrieval setting, despite variations in UAV altitudes and satellite image resolutions, our method effectively captures spatial-geographic cues, achieving high retrieval accuracy while substantially reducing mismatches.
-Compared to University-1652, DenseUAV provides a more realistic and complex benchmark for UAV-based geo-localization.
-Under this setting, our method exhibits strong cross-view retrieval performance, demonstrating robust feature discrimination in densely distributed urban regions.
+In the University-1652 dataset, the dronesatellite retrieval task uses drone-view images as queries to retrieve corresponding satellite images.
+Our method achieves consistent Top-1 accuracy to highlight its effectiveness in cross-view matching.
+In the satellitedrone setting, our method also accurately retrieves drone-view images from satellite queries that demonstrates strong cross-view retrieval ability.
+Compared with University-1652, the DenseUAV dataset introduces more realistic and complex challenges, such as dense spatial sampling, multi-scale resolutions, and temporal dynamics.
+In the drone satellite retrieval setting, despite the high risk of confusion caused by dense spatial sampling, our method effectively captures spatial-geographic cues and achieves high retrieval accuracy, demonstrating robust cross-view performance under realistic and complex conditions.
 V CONCLUSION
-This paper investigates a highly valuable and challenging task namely self-supervised UAV-view cross-view geo-localization, aiming to reduce the reliance on paired UAV–satellite images.
-To this end, we propose a self-supervised UVGL method, which employs a shallow backbone with only 9 layers for feature extraction.
-Pseudo labels are generated via clustering, and a dual-branch contrastive learning mechanism is introduced to enhance the capability in representing intra-view visual patterns.
+This paper investigates a highly valuable and challenging DVGL task to reduce reliance on paired drone–satellite images.
+To this end, we propose a self-supervised learning method that employs a shallow backbone with only 9 layers for feature extraction.
+Pseudo-labels are generated via clustering, and a dual-branch contrastive learning mechanism is introduced to enhance the capability in representing intra-view visual patterns.
 To further improve intra-view discriminability and consistency, we design a dynamic hierarchical memory learning module.
-Additionally, to mitigate the structural and semantic domain gap between views, we propose an information consistency evolution learning module, which facilitates semantic alignment by learning high-dimensional feature consistency across views.
-To improve the quality and robustness of pseudo labels, we introduce a bidirectional feature consistency-based Pseudo Label Enhancement strategy, which refines pseudo-supervision signals and boosts cross-view representation learning.
+Additionally, to mitigate the structural and semantic domain gap between views, we propose an information consistency evolution learning module, which facilitates semantic alignment by learning high-dimensional feature consistency across-views.
+To improve the quality and robustness of pseudo-labels, we introduce a pseudo-label enhancement strategy, which refines pseudo-supervision signals and boosts cross-view representation learning.
 As a result, the proposed method effectively unifies both intra-view and cross-view feature representations in the embedding space.
-Extensive experiments conducted on three public benchmarks University-1652, SUES-200, and DenseUAV, demonstrate that our method significantly outperforms existing self-supervised methods, even surpassing several fully supervised methods in certain scenarios, without relying on paired UAV–satellite images.
-These results validate the effectiveness and generalization capability of our method, highlighting its potential for real-world UVGL applications.
+Extensive experiments conducted on three public benchmarks University-1652, SUES-200, and DenseUAV, demonstrate that our method significantly outperforms existing self-supervised methods, and even surpasses several fully supervised methods in certain scenarios without relying on paired drone–satellite images.
+These results validate the effectiveness and generalization capability of our method and verify the potential for real-world DVGL applications.